### v1

# Parameter-Efficient Fine-Tuning via Circular Convolution

Abstract
Low-Rank Adaptation (LoRA) has gained popularity for fine-tuning large foundation models, leveraging low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ to represent weight changes (i.e., $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$). This method reduces trainable parameters and mitigates heavy memory consumption associated with full delta matrices by sequentially multiplying $\mathbf{A}$ and $\mathbf{B}$ with the activation. Despite its success, the intrinsic low-rank characteristic may limit its performance. Although several variants have been proposed to address this issue, they often overlook the crucial computational and memory efficiency brought by LoRA. In this paper, we propose Circular Convolution Adaptation (C3A), which not only achieves high-rank adaptation with enhanced performance but also excels in both computational power and memory utilization. Extensive experiments demonstrate that C3A consistently outperforms LoRA and its variants across various fine-tuning tasks.
1 Introduction
In recent years, Large Foundation Models (LFMs) have surged in popularity due to their remarkable performance across various tasks in natural language processing Brown et al. (2020); Touvron et al. (2023), computer vision Radford et al. (2021); Kirillov et al. (2023), and other domains Li et al. (2024). Characterized by their vast number of parameters and substantial computational demands, these models have set new benchmarks in accuracy and versatility. However, their size and complexity pose significant challenges for efficient fine-tuning, particularly in resource-constrained environments Malladi et al. (2023); Zhang et al. (2024). To address this issue, Parameter-efficient fine-tuning (PEFT) techniques Mangrulkar et al. (2022), such as Low-Rank Adaptation (LoRA) Hu et al. (2021), have emerged as practical solutions.
LoRA reduces the number of trainable parameters by employing low-rank matrices to approximate weight changes, making fine-tuning more feasible and accessible without compromising the model’s performance. Concretely, LoRA can be expressed in math:
$\mathbf{Wx}=(\mathbf{W}_{0}+\Delta\mathbf{W})\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B(Ax)},$
where $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d\times d}$ are weight matrices, $\mathbf{B}\in\mathbb{R}^{d\times r},\mathbf{A}\in\mathbb{R}^{r\times d}$ are low-rank matrices to construct $\Delta\mathbf{W}$ and $\mathbf{x}\in\mathbb{R}^{d}$ is the activations. The number of trainable parameters is $2rd$, prompting the choice $r\ll d$ (e.g., $r=8$ for $d=1024$) for achieving high parameter efficiency. However, as discussed in Zeng & Lee (2023), LoRA’s capacity to represent a target model is strictly limited by $r$. To address this trade-off between performance and efficiency, Kopiczko et al. (2023) proposed Vector Random Matrix Adaptation (VeRA), which achieves competitive performance with significantly fewer trainable parameters through fixed random-matrix projections. Despite its low parameter count, VeRA requires substantial computational power and memory due to the large size of the random matrices used in projection. This raises the following open research question in the context of PEFT:
How can high-rank adaptation be achieved without incurring significant costs in terms of computational power, memory usage, and trainable parameters?
As discussed in Dosovitskiy et al. (2020), dense linear layers lack strong inductive biases, limiting transformers’ ability to compete with Convolutional Neural Networks (CNNs) when data is scarce. Given the constraints of a small training set for the downstream task, we posit that incorporating a stronger inductive bias could enhance adaptation. In light of this, we propose Circular Convolution Adaptation (C3A). Circular convolution is famous in signal processing Li et al. (2020) and cryptography Dworkin et al. (2001). This operator can be equivalently represented as multiplication by a circulant matrix, offering flexibility in rank unconstrained by the number of trainable parameters. Moreover, leveraging the Fast Fourier Transform (FFT), C3A ensures better computational and memory efficiency than directly multiplying a delta matrix Bamieh (2018), rendering it competitive with LoRA from the efficiency aspect. We summarize our contribution as below:
① We introduce C3A, a novel approach for PEFT using convolution. This method leverages circular convolution kernels and their equivalent circulant matrices to provide a flexible rank, which is free of linear constraint by the number of trainable parameters, for the delta matrix.
② Leveraging the elegant diagonalization of the circulant matrix, we implement both the forward pass and backpropagation using FFT. With the incorporation of FFT, the computation and memory efficiency of C3A excels. C3A strikes a unique balance between performance an efficiency.
③ To offer greater flexibility in controlling the number of trainable parameters, we extend C3A by incorporating block-circular convolution, which results in block-circulant matrices. This extension allows C3A to achieve fully customizable parameter counts as well as adaptable rank configurations.
④ We validate C3A through comprehensive fine-tuning experiments across diverse tasks including natural language understanding, instruction tuning, image classification, and text-to-image generation. Our experiments demonstrate C3A’s outstanding accuracy and memory merits compared to existing state-of-the-art methods.
2 Related Work
2.1 Parameter-Efficient Fine-Tuning
Research on Parameter-Efficient Fine-Tuning (PEFT) has generally progressed along three main directions. The first direction involves partially updating the pre-trained neural network (e.g., the body Oh et al. (2020) or the biases Zaken et al. (2021)). Initial methods relied on hand-crafted heuristics Han et al. (2015); Oh et al. (2020) to identify which parameters are crucial and should be fine-tuned. More advanced approaches employ optimization techniques Guo et al. (2020); Xu et al. (2021); Fu et al. (2023). For example, Guo et al. Guo et al. (2020) reformulated the discrete optimization problem into a continuous one using Bernoulli masks and the Gumbel-softmax approximation.
The second direction emerged to maintain the integrity of the pre-trained model and enable a high degree of parameter sharing through adapter-based methods He et al. (2021); Rebuffi et al. (2017); Rücklé et al. (2020); Liu et al. (2022). These works focus on integrating additional modules, termed adapters, to fit the downstream task, effectively decoupling the pre-trained model parameters from those specific to the downstream task. Prompt Tuning Brown et al. (2020); Gao et al. (2020); Chen et al. (2023) and Prefix Tuning Li & Liang (2021) also fall into this category, despite ignoring potential semantic meanings.
The final direction is characterized by delta-weight-based methods, such as Low-Rank Adaptation (LoRA) Hu et al. (2021) and Orthogonal Fine-tuning (OFT) Qiu et al. (2023). These methods bridge the gap between the pre-trained model and the downstream task by adaptive delta weights, which is learned and stored separately while used in combination with the pre-trained weights. This unique design enables disentanglement and parameter sharing while preserving the ability of integrating without additional inference cost. LoRA models the delta-weights by an additive matrix while OFT does it by a multiplicative one. To further improve either parameter efficiency or performance, many variants has been proposed for both of the methods Kopiczko et al. (2023); Liu et al. (2024; 2023); Yuan et al. (2024). However, these methods can hardly achieve high parameter efficiency and performance without incurring heavy computation and memory usage.
2.2 Circular Convolution
Circular convolution has been extensively studied in signal processing Rabiner et al. (1978); McGillem & Cooper (1984); Li et al. (2020) and cryptography Dworkin et al. (2001); Gong et al. (2024). Owing to its computational advantages, circular convolution has also been explored in machine learning for generating long embeddings of high-dimensional data Yu et al. (2014) and compressing heavily parameterized layers Cheng et al. (2015); Ding et al. (2017). Remarkably, it achieves these efficiencies without significant performance degradation, which makes it a promising technique for fine-tuning applications.
Despite its success in small neural networks such as LeNet Cheng et al. (2015), circular convolution has not demonstrated lossless performance in modern large foundational models (LFMs) or even in their base architecture, the transformer. This limitation may be attributed to the conflict between its high intrinsic bias (i.e., the circulant pattern) and the vast amount of data required for training LFMs. Conversely, when fine-tuning LFMs, it is often impractical to collect as much data as needed for training from scratch. In such scenarios, the intrinsic bias of circular convolution could potentially serve as a regularization mechanism, thereby benefiting the fine-tuning process.
3 Method
In this section, we present C3A (see an overview in Figure 1), a novel PEFT approach based on the circular convolution. C3A follows LoRA’s setting of learning an additive linear operation over the original dense linear transformation. However, instead of using low-rank decomposition and the matrix multiplication operator, C3A resorts to circular convolution as this additive linear operation. Section 3.1 introduces the notations we use. Section 3.2 discusses the circular convolution operator, its equivalent circulant matrix, and its calculation in the frequency domain. Section 3.3 details an efficient method for backpropagation. Section 3.4 describes block-wise convolution for controlling the number of trainable parameters. Finally, Section 3.5 analyzes the computational complexity.
Figure 1: Overview of LoRA (A) and our C3A (B,C) method. In LoRA, only low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ are trained and the delta weight is represented by their product (i.e., $\Delta\mathbf{W}=\mathbf{BA}$). The total trainable parameter number is $2rd$, which is assosiated with the rank of the delta weight. In C3A, circular convolution kernels $\mathbf{w}$ are tuned to adapt to the downstream task and the delta weight is represented by the (block-)circular matrix they construct (i.e., $\Delta\mathbf{W}=\mathcal{C}_{b}(\Delta\mathbf{w})$). The total trainable parameter number is $bd$, which disentangles with the rank of the delta weight.
3.1 Notations
We denote circular convolution operator as $\star$. The adapted weight matrix, the original weight matrix, and the delta matrix are denoted by $\mathbf{W}$, $\mathbf{W}_{0}$, and $\Delta\mathbf{W}$, respectively. The activation vector of the previous layer is denoted by $\mathbf{x}$. The post-transformation vector is $\mathbf{z}$, where $\mathbf{z}=\mathbf{W}\mathbf{x}$, and the incremental part is denoted by $\Delta\mathbf{z}$, where $\Delta\mathbf{z}=\Delta\mathbf{W}\mathbf{x}$. The length of the activation vector is $d$. The matrices $\mathbf{A}$ and $\mathbf{B}$ are low-rank matrices introduced by LoRA to represent $\Delta\mathbf{W}$, with $r$ being their rank. $r_{v}$ is used to specify the random projection matrix rank used in VeRA. The circular convolution kernel of C3A is denoted by $\Delta\mathbf{w}$. The loss function is represented by $\mathcal{L}$. The Fast Fourier Transform and its inverse are denoted by $\mathrm{FFT}$ and $\mathrm{iFFT}$, respectively. The Hadamard product is denoted by $\circ$. In terms of dimensions, we have $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d\times d}$, and $\mathbf{x},\mathbf{z},\Delta\mathbf{z},\Delta\mathbf{w}\in\mathbb{R}^{d}$. The matrices $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d}$ are controlled by the rank $r$.
3.2 Circular Convolution
The circular convolution operator is defined as $\Delta\mathbf{z}=\Delta\mathbf{w}\star\mathbf{x}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, where $\mathcal{C}(\Delta\mathbf{w})$ is the circulant matrix of $\Delta\mathbf{w}$. The first row of $\mathcal{C}(\Delta\mathbf{w})$ is $\Delta\mathbf{w}$ and each row equals the row above it periodically shifted to right by one element. Concretely,
$\mathcal{C}(\Delta\mathbf{w})=\begin{bmatrix}\Delta w_{1}&\Delta w_{2}&\cdots&\Delta w_{d-1}&\Delta w_{d}\\ \Delta w_{d}&\Delta w_{1}&\cdots&\Delta w_{d-2}&\Delta w_{d-1}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ \Delta w_{3}&\Delta w_{4}&\cdots&\Delta w_{1}&\Delta w_{2}\\ \Delta w_{2}&\Delta w_{3}&\cdots&\Delta w_{d}&\Delta w_{1}\\ \end{bmatrix}\in\mathbb{R}^{d\times d}.$
Theoretically, the rank of $\mathcal{C}(\Delta\mathbf{w})$ is $d-\mathrm{Deg}(\mathrm{gcd}(f(x),x^{d}-1))$ Ingleton (1956), where $\mathrm{Deg}(\cdot)$ represents the degree of a polynomial, $f(x)$ is the associated polynomial of $\Delta\mathbf{w}$ (i.e., $f(x)=\sum_{i=1}^{d}\Delta w_{i}x^{i-1}$), and $\mathrm{gcd}(\cdot)$ means the greatest common divisor. By learning $\Delta\mathbf{w}$ in the $\mathbb{R}^{n}$ oracle, C3A automatically achieves dynamic rank selection that is not linearly bounded by the number of learnable parameters as LoRA.
To achieve efficiency, we leverage the beautiful circulant structure of $\mathcal{C}(\Delta\mathbf{w})$, which makes it diagonalizable by the Fourier basis ($\mathbf{F}$) . In math, it can be described as $\mathcal{C}(\Delta\mathbf{w})=\mathbf{F}\frac{\Lambda}{d}\mathbf{F}^{-1}$ Golub & Van Loan (1996), where $\Lambda$ is its eigenvalues and can be calculated by a Fourier transform of the first row (i.e., $\Lambda=\mathrm{diag}(\mathbf{F}\Delta\mathbf{w})$). Therefore, we can calculate $\Delta\mathbf{w}\star\mathbf{x}$ as
$\begin{split}\Delta\mathbf{w}\star\mathbf{x}&=\mathbf{F}\mathrm{diag}(\frac{\mathbf{F}\Delta\mathbf{w}}{d})\mathbf{F}^{-1}\mathbf{x}\\ &=\mathrm{FFT}(\mathrm{FFT}(\Delta\mathbf{w})\circ\mathrm{iFFT}(\mathbf{x})).\end{split}$ (1)
3.3 Backpropagation
To implement backpropagation efficiently, we require the analytical derivatives of the loss function $\mathcal{L}$, with respect to $\Delta\mathbf{w}$ and $\mathbf{x}$. Utilizing the chain rule, these derivatives can be expressed as:
$\frac{\partial\mathcal{L}}{\partial\mathbf{x}}=\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}},\qquad\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{w}}=\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}}.$ (2)
Since $\Delta\mathbf{z}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, it follows that $\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})$. For $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}$, we note that the circular convolution operation is commutative (i.e., $\Delta\mathbf{w}\star\mathbf{x}=\mathbf{x}\star\Delta\mathbf{w}$), thus $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})$. Substituting these results into Equation 2, we obtain:
$\frac{\partial\mathcal{L}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}},\qquad\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}}.$
These expressions can also be interpreted as circular convolutions:
$\frac{\partial\mathcal{L}}{\partial\mathbf{x}}=\Delta\mathbf{w}\star\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}},\qquad\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{w}}=\mathbf{x}\star\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}}.$
By explicitly implementing this derivative calculation using Equation 1, backpropagation can also enjoy the computational efficiency provided by the FFT algorithm.
3.4 Block-Circular Convolution
Despite the elegance and efficiency of the circular convolution operator, it is worth noting that we can not alter the number of learnable parameters due to the requirement of the convolution kernel being the same size as the activation vector. To enable this flexibility in case of difficult downstream tasks, we resort to the block-circular convolution. We can simply split the activation vector $\mathbf{x}$ (and the post-transformation vector $\mathbf{z}$) into blocks, and assign each pair of blocks with different convolution kernels. To be concrete,
$\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\cdots&\mathbf{x}_{b}\end{bmatrix}\qquad\Delta\mathbf{z}=\begin{bmatrix}\Delta\mathbf{z}_{1}&\Delta\mathbf{z}_{2}&\cdots&\Delta\mathbf{z}_{b}\end{bmatrix},$
where $b$ is the block number. We will need $b^{2}$ convolution kernels to densely connect these blocks, which can be expressed as
$\Delta\mathbf{z}_{i}=\sum_{j=1}^{b}\Delta\mathbf{w}_{ij}\star\mathbf{x}_{j},i\in\{1,2,\cdots,b\},$
and can also be represented by a block-circular matrix
$\Delta\mathbf{z}=\mathcal{C}_{b}(\Delta\mathbf{w})\mathbf{x}\qquad\mathcal{C}_{b}(\Delta\mathbf{w})=\begin{bmatrix}\mathcal{C}(\Delta\mathbf{w}_{11})&\mathcal{C}(\Delta\mathbf{w}_{12})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{1b})\\ \mathcal{C}(\Delta\mathbf{w}_{21})&\mathcal{C}(\Delta\mathbf{w}_{22})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{2b})\\ \cdots&\cdots&\cdots&\cdots\\ \mathcal{C}(\Delta\mathbf{w}_{b1})&\mathcal{C}(\Delta\mathbf{w}_{b2})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{bb})\end{bmatrix}.$
In this case, $\Delta\mathbf{w}_{ij}\in\mathbb{R}^{\frac{d}{b}}$ (it is required that $d$ is divisible by $b$) and we have $b^{2}\frac{d}{b}=bd$ learnable parameters. $b$ acts as a hyperparameter controlling the number of learnable parameters here, just like $r$ for LoRA. However, we highlight that while $r$ controls the rank of the delta matrix and learnable parameter number at the same time, $b$ only affects the latter. Disentangling these attributes enables higher adaptation flexibility and potential better results.
3.5 Complexity Analysis
We compare time compelxity and space complexity of LoRA, VeRA and C3A in Table 1. Detailed analysis follows.
Table 1: Time and space complexity comparison of LoRA, VeRA and C3A. We split the space complexity into Parameter number and Other auxiliary tensors to help better understand the differences. We highlight that in practice, to achieve similar performance, $b\leq r\ll r_{v}$.
Method Time # Parameter # Other # Total
LoRA $\mathcal{O}(rd)$ $2rd$ $r$ $2rd+r$
VeRA $\mathcal{O}(r_{v}d)$ $r_{v}+d$ $2r_{v}d+r_{v}+d$ $2(r_{v}d+r_{v}+d)$
C3A $\mathcal{O}(bd\log\frac{d}{b})$ $bd$ 00 $bd$
3.5.1 Time Complexity
LoRA incorporates low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, which are sequentially multiplied with the activation vector, resulting in a time complexity of $\mathcal{O}(rd)$. Typically, $r\ll d$. Due to the $\mathcal{O}(d\log d)$ complexity of the FFT algorithm used in Equation 1, C3A achieves a time complexity of $\mathcal{O}(bd\log\frac{d}{b})$. Given that $b\ll d$, C3A is comparable to LoRA in terms of time complexity. Conversely, VeRA, despite its high-rank nature and minimal number of trainable parameters, incurs a prohibitive time complexity of $\mathcal{O}(r_{v}d)$, where $r_{v}$ can exceed $d$. Therefore, striking a balance between high rank and computational efficiency is never a low-hanging fruit.
3.5.2 Space Complexity
We analyze the space complexity of LoRA, VeRA, and C3A during training. The differences among these methods primarily arise from the trainable parameter size and the auxiliary tensors required for the forward pass and backpropagation. For LoRA, an intermediate tensor of size $r$ is introduced (i.e., $\mathbf{Ax}\in\mathbb{R}^{r}$). In contrast, VeRA necessitates two random projection matrices, with a total size of $2r_{v}d$. Additionally, VeRA introduces two intermediate tensors of sizes $d$ and $r_{v}$. Thanks to the efficient backpropagation algorithm detailed in Section 3.3, C3A does not require any additional intermediate tensors. This optimization significantly reduces its space complexity compared to LoRA and VeRA.
4 Experiment
In this section, we comprehensively evaluate C3A in both natural language processing (NLP) and computer vision (CV). For NLP, we first demonstrate the effectiveness of C3A using the GLUE benchmark with RoBERTa-Base and RoBERTa-Large models. Additionally, we apply C3A to fine-tuning the LLaMA family models, which stands as a representative of prevalent large language models (LLMs). For CV, we examine two practical scenarios: the traditional classification task using Vision Transformers (ViTs) and the more intricate text-to-image generation task based on diffusion models.
4.1 Natural Language Understanding
Baselines. We compare our C3A with several representative PEFT methods, including BitFit Zaken et al. (2021), (IA)3 Liu et al. (2022), LoRA Hu et al. (2021), VeRA Kopiczko et al. (2023), and BOFT Liu et al. (2023). BitFit selectively fine-tunes existing parameters, specifically the biases. (IA)3 is the state-of-the-art method that adds extra adapters. LoRA is a widely known PEFT method that employs low-rank decomposition to compress additive delta matrices. VeRA is a recent approach that focuses on further reducing trainable parameters for LoRA. BOFT is another innovative method in PEFT research, compressing multiplicative delta matrices using orthogonal decomposition and butterfly factorization.
Settings. We evaluate our proposed C3A on the General Language Understanding Evaluation (GLUE) benchmark Wang et al. (2018), which encompasses a wide range of natural language understanding (NLU) tasks, including single-sentence classification, similarity and paraphrase, and natural language inference. To enhance practicality, we split these datasets following the train-validation-test approach. The best-performing model is selected based on validation set performance across the fine-tuning epochs, and the reported performance corresponds to its performance on the test set. For this evaluation, we fine-tune the pre-trained RoBERTa-Base and RoBERTa-Large models Liu et al. (2019). For the unique hyperparameters of different baselines, we adopt the values suggested in the original papers (e.g., VeRA’s $r$ and BOFT’s $b$ and $m$). The number of trainable parameters excludes the classification head, as each method uses the same one. The shared hyperparameters (i.e., the learning rate for classification head and the learning rate for other trainable parameters) are found by hyperparameter search. For the memory cost, to ensure fairness and consistency, we fix the length of input data to 256 tokens and use a batchsize of 64.
Results. Results are summarized in Table 2. In general, C3Ab=1 achieves better or on-par performance compared to baseline methods with extremely small trainable parameters. When a larger number of trainable parameters is allowed, C3Ab=6 and C3Ab=8 significantly outperform the baselines. Furthermore, compared to (IA)3, LoRA, VeRA, and BOFT, despite training a similar number of or more parameters, C3A stands out for its excellent memory efficiency. The only method with better memory efficiency is BitFit, which can be seen as an upper bound since it does not introduce any new parameters.
Table 2: Performance of different PEFT methods on the GLUE benchmark. We fine-tune pre-trained RoBERTa-Base and -Large models on 6 datasets. We report the Matthew’s correlation coefficient (MCC) for CoLA, Pearson correlation coefficient (PCC) for STS-B, and accuracy (Acc.) for all remaining tasks. For all 6 datasets, a higher metric indicates better performance. “Avg.” denotes the average accuracy of each method across all datasets. The best results for each dataset are highlighted in bold. # Trainable parameters does not include the classification head since each method uses the same head. Memory Cost is measured on fixed length (i.e., $256$) data with a batchsize of $64$.
Methods # Trainable Parameters Memory Cost (GB) SST-2 MRPC CoLA QNLI RTE STS-B Avg.
Base BitFit 102K 12.60 93.30$\pm$0.30 85.80$\pm$0.21 59.21$\pm$1.74 91.96$\pm$0.18 73.07$\pm$1.34 90.18$\pm$0.17 82.25
(IA)3  [^1] K 19.86 92.98$\pm$0.34 85.86$\pm$0.59 60.49$\pm$1.09 91.56$\pm$0.17 69.10$\pm$1.18 90.06$\pm$0.21 81.67
LoRAr=8 295K 13.75 94.50$\pm$0.41 85.68$\pm$0.74 60.95$\pm$1.57 92.54$\pm$0.20 76.68$\pm$1.42 89.76$\pm$0.39 83.35
VeRAr=1024 43K 15.51 93.97$\pm$0.17 86.23$\pm$0.41 62.24$\pm$1.91 91.85$\pm$0.17 75.74$\pm$1.56 90.27$\pm$0.25 83.38
[…]
C3Ab=1 18K 12.83 93.42$\pm$0.26 86.33$\pm$0.32 61.83$\pm$0.96 91.83$\pm$0.04 76.17$\pm$1.39 90.46$\pm$0.29 83.34
C3Ab=6  [^1] K 12.72 94.20$\pm$0.16 86.67$\pm$0.54 62.48$\pm$1.20 92.32$\pm$0.25 77.18$\pm$1.41 90.16$\pm$0.42 83.84
Large BitFit 271K 30.65 95.09$\pm$0.27 88.10$\pm$0.76 65.40$\pm$0.76 94.06$\pm$0.14 82.60$\pm$1.15 91.73$\pm$0.20 86.16
(IA)3 295K 48.81 95.32$\pm$0.20 87.06$\pm$0.57 66.52$\pm$1.10 94.18$\pm$0.15 84.33$\pm$2.38 91.58$\pm$0.39 86.50
LoRAr=8 786K 34.12 95.53$\pm$0.35 86.12$\pm$3.22 65.16$\pm$0.76 93.73$\pm$0.30 83.75$\pm$0.51 91.46$\pm$0.21 85.96
VeRAr=256 61K 34.16 95.83$\pm$0.43 87.72$\pm$0.55 63.66$\pm$1.45 94.11$\pm$0.20 83.03$\pm$1.65 91.12$\pm$0.37 85.91
[…]
C3Ab=1 49K 31.83 95.78$\pm$0.05 88.02$\pm$0.62 66.59$\pm$1.20 94.22$\pm$0.25 82.89$\pm$0.67 91.86$\pm$0.14 86.56
C3Ab=8 393K 31.79 95.78$\pm$0.15 88.09$\pm$0.47 67.18$\pm$1.92 94.26$\pm$0.19 84.62$\pm$1.36 91.81$\pm$0.36 86.96
5 Conclusion
In this paper, we introduce C3A, an innovative approach to Parameter-Efficient Fine-Tuning (PEFT). Unlike LoRA, which utilizes low-rank decomposition, C3A employs circular convolution and its associated circulant matrix to model the delta weight matrix. This approach aims to decouple the rank of the delta weight matrix from the number of learnable parameters, allowing for high-rank adaptation while maintaining a limited parameter size. By leveraging the Fast Fourier Transform (FFT) in both the forward and backward passes, C3A achieves significant computational and memory efficiency. Consequently, C3A presents itself as a compelling alternative to LoRA for model fine-tuning.
Appendix A Appendix

### v2

@@
 1 Introduction
-In recent years, Large Foundation Models (LFMs) have surged in popularity due to their remarkable performance across various tasks in natural language processing Brown et al.
+Figure 1: Relative comparison of C3A and baselines on RoBERTa-Base.
+The Pearson Correlation Coefficient (PCC) is evaluated on STS-B and the Matthew’s Correlation Coefficient (MCC) on CoLA.
+Accuracies across SST-2, MRPC, QNLI, and RTE are averaged and reported as Acc.
+#Param denotes the count of learnable parameters (excluding the head), while Memory indicates the peak GPU memory usage during training.
+Metrics highlighted in blue represent values where higher is better, and those in red indicate values where lower is better.
+Refer to Table 2 for absolute statistics.
+In recent years, Large Foundation Models (LFMs) have surged in popularity due to their remarkable performance across various tasks in natural language processing (NLP) Brown et al.
 (2020); Touvron et al.
-(2023), computer vision Radford et al.
+(2023), computer vision (CV) Radford et al.
 (2021); Kirillov et al.
@@
 (2023); Zhang et al.
-(2024).
+(2024b).
 To address this issue, Parameter-efficient fine-tuning (PEFT) techniques Mangrulkar et al.
 (2022), such as Low-Rank Adaptation (LoRA) Hu et al.
-(2021), have emerged as practical solutions.
+(2021), have emerged as outstanding practical solutions.
 LoRA reduces the number of trainable parameters by employing low-rank matrices to approximate weight changes, making fine-tuning more feasible and accessible without compromising the model’s performance.
@@
 $\mathbf{Wx}=(\mathbf{W}_{0}+\Delta\mathbf{W})\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B(Ax)},$
-where $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d\times d}$ are weight matrices, $\mathbf{B}\in\mathbb{R}^{d\times r},\mathbf{A}\in\mathbb{R}^{r\times d}$ are low-rank matrices to construct $\Delta\mathbf{W}$ and $\mathbf{x}\in\mathbb{R}^{d}$ is the activations.
-The number of trainable parameters is $2rd$, prompting the choice $r\ll d$ (e.g., $r=8$ for $d=1024$) for achieving high parameter efficiency.
+where $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ are weight matrices, $\mathbf{B}\in\mathbb{R}^{d_{1}\times r},\mathbf{A}\in\mathbb{R}^{r\times d_{2}}$ are low-rank matrices to construct $\Delta\mathbf{W}$ and $\mathbf{x}\in\mathbb{R}^{d_{2}}$ is the activations.
+The number of trainable parameters is $r(d_{1}+d_{2})$, prompting the choice $r\ll\min(d_{1},d_{2})$ (e.g., $r=8$ for $d_{1}=d_{2}=1024$) for achieving high parameter efficiency.
 However, as discussed in Zeng & Lee (2023), LoRA’s capacity to represent a target model is strictly limited by $r$.
@@
 (2023) proposed Vector Random Matrix Adaptation (VeRA), which achieves competitive performance with significantly fewer trainable parameters through fixed random-matrix projections.
-Despite its low parameter count, VeRA requires substantial computational power and memory due to the large size of the random matrices used in projection.
+Despite its low parameter count, VeRA requires substantial computational power and memory due to the large size of the random matrices used for projection.
 This raises the following open research question in the context of PEFT:
-How can high-rank adaptation be achieved without incurring significant costs in terms of computational power, memory usage, and trainable parameters?
+How can high-rank adaptation be achieved without incurring significant costs in terms of time complexity, memory usage, and trainable parameters?
 As discussed in Dosovitskiy et al.
-(2020), dense linear layers lack strong inductive biases, limiting transformers’ ability to compete with Convolutional Neural Networks (CNNs) when data is scarce.
-Given the constraints of a small training set for the downstream task, we posit that incorporating a stronger inductive bias could enhance adaptation.
-In light of this, we propose Circular Convolution Adaptation (C3A).
-Circular convolution is famous in signal processing Li et al.
+(2020), dense linear layers lack inductive biases, which poses challenge on the optimization process and therefore limits transformers’ ability to compete with Convolutional Neural Networks (CNNs) when data is scarce.
+Given the constraints of a small training set for the downstream task, we posit that incorporating a stronger inductive bias could improve the adaptation performance.
+In light of this, we propose Circular Convolution Adaptation (C3A), featuring the circular convolution operator Bamieh (2018).
+Figure 2: Overview of LoRA (A) and our C3A (B,C) method.
+In LoRA, only low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ are trained and the delta weight is represented by their product (i.e., $\Delta\mathbf{W}=\mathbf{BA}$).
+The total trainable parameter number is $r(d_{1}+d_{2})$, which is assosiated with the rank of the delta weight.
+In C3A, circular convolution kernels $\Delta\mathbf{w}$ are tuned to adapt to the downstream task and the delta weight is represented by the (block-)circular matrix they construct (i.e., $\Delta\mathbf{W}=\mathcal{C}_{\mathrm{(blk)}}(\Delta\mathbf{w})$).
+The total trainable parameter count is $\frac{d_{1}d_{2}}{b}$, which disentangles with the rank of the delta weight.
+Here, $b$ is the block size of the block-circular matrix and it should be a common divisor (CD) of $d_{1}$ and $d_{2}$.
+Circular convolution has been famous in signal processing Li et al.
 (2020) and cryptography Dworkin et al.
-(2001).
-This operator can be equivalently represented as multiplication by a circulant matrix, offering flexibility in rank unconstrained by the number of trainable parameters.
-Moreover, leveraging the Fast Fourier Transform (FFT), C3A ensures better computational and memory efficiency than directly multiplying a delta matrix Bamieh (2018), rendering it competitive with LoRA from the efficiency aspect.
-We summarize our contribution as below:
+(2001) for its distinguished efficiency and compactness.
+This operator can be equivalently represented as multiplication by a circulant matrix, which offers rank flexibility that is unconstrained by the number of trainable parameters.
+Moreover, leveraging the Fast Fourier Transform (FFT), C3A ensures better time and memory efficiency than directly multiplying the circulant matrix Bamieh (2018), rendering it competitive with LoRA from the efficiency aspect.
+In a word, circular convolution is a promising option for escaping the rank curse of LoRA at minimal cost.
+We summarize our contribution as follows:
 ① We introduce C3A, a novel approach for PEFT using convolution.
-This method leverages circular convolution kernels and their equivalent circulant matrices to provide a flexible rank, which is free of linear constraint by the number of trainable parameters, for the delta matrix.
+This method leverages the circular convolution operation and its equivalent circulant matrix to provide a flexible rank, which is free of linear constraint by the number of trainable parameters, for the delta matrix.
 ② Leveraging the elegant diagonalization of the circulant matrix, we implement both the forward pass and backpropagation using FFT.
@@
 This extension allows C3A to achieve fully customizable parameter counts as well as adaptable rank configurations.
-④ We validate C3A through comprehensive fine-tuning experiments across diverse tasks including natural language understanding, instruction tuning, image classification, and text-to-image generation.
-Our experiments demonstrate C3A’s outstanding accuracy and memory merits compared to existing state-of-the-art methods.
+④ We validate C3A through comprehensive fine-tuning experiments across diverse tasks including natural language understanding, instruction tuning and image classification.
+Experiments demonstrate C3A’s outstanding accuracy and memory merits compared to existing state-of-the-art methods.
 2 Related Work
@@
 (2021)).
-Initial methods relied on hand-crafted heuristics Han et al.
+Traditional methods relied on hand-crafted heuristics Han et al.
 (2015); Oh et al.
@@
 For example, Guo et al.
-Guo et al.
-(2020) reformulated the discrete optimization problem into a continuous one using Bernoulli masks and the Gumbel-softmax approximation.
-The second direction emerged to maintain the integrity of the pre-trained model and enable a high degree of parameter sharing through adapter-based methods He et al.
+(2020) reformulated such a discrete optimization problem into a continuous one by employing Bernoulli masks and the Gumbel-softmax approximation.
+The second direction emerged to maintain the integrity of the pre-trained model while enabling a high degree of parameter sharing through adapter-based methods He et al.
 (2021); Rebuffi et al.
@@
 (2020); Chen et al.
-(2023) and Prefix Tuning Li & Liang (2021) also fall into this category, despite ignoring potential semantic meanings.
+(2023); Zhang et al.
+(2024a) and Prefix Tuning Li & Liang (2021) also fall into this category, despite ignoring potential semantic meanings.
 The final direction is characterized by delta-weight-based methods, such as Low-Rank Adaptation (LoRA) Hu et al.
@@
 (2023).
-These methods bridge the gap between the pre-trained model and the downstream task by adaptive delta weights, which is learned and stored separately while used in combination with the pre-trained weights.
-This unique design enables disentanglement and parameter sharing while preserving the ability of integrating without additional inference cost.
+These methods bridge the gap between the pre-trained model and the downstream task by adaptive delta weights, which are stored separately while used in combination with the pre-trained weights.
+This unique design enables disentanglement of the pretrained and downstream-specific weights.
+Namely, it achieves parameter sharing and preserves the ability to integrate without additional inference cost.
 LoRA models the delta-weights by an additive matrix while OFT does it by a multiplicative one.
@@
 Conversely, when fine-tuning LFMs, it is often impractical to collect as much data as needed for training from scratch.
-In such scenarios, the intrinsic bias of circular convolution could potentially serve as a regularization mechanism, thereby benefiting the fine-tuning process.
+In such scenarios, the intrinsic bias of circular convolution could potentially serve as a regularization mechanism, thereby benefiting the optimization process of fine-tuning.
 3 Method
@@
 Finally, Section 3.5 analyzes the computational complexity.
-Figure 1: Overview of LoRA (A) and our C3A (B,C) method.
-In LoRA, only low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ are trained and the delta weight is represented by their product (i.e., $\Delta\mathbf{W}=\mathbf{BA}$).
-The total trainable parameter number is $2rd$, which is assosiated with the rank of the delta weight.
-In C3A, circular convolution kernels $\mathbf{w}$ are tuned to adapt to the downstream task and the delta weight is represented by the (block-)circular matrix they construct (i.e., $\Delta\mathbf{W}=\mathcal{C}_{b}(\Delta\mathbf{w})$).
-The total trainable parameter number is $bd$, which disentangles with the rank of the delta weight.
 3.1 Notations
-We denote circular convolution operator as $\star$.
-The adapted weight matrix, the original weight matrix, and the delta matrix are denoted by $\mathbf{W}$, $\mathbf{W}_{0}$, and $\Delta\mathbf{W}$, respectively.
-The activation vector of the previous layer is denoted by $\mathbf{x}$.
-The post-transformation vector is $\mathbf{z}$, where $\mathbf{z}=\mathbf{W}\mathbf{x}$, and the incremental part is denoted by $\Delta\mathbf{z}$, where $\Delta\mathbf{z}=\Delta\mathbf{W}\mathbf{x}$.
-The length of the activation vector is $d$.
+The adapted weight matrix, the original weight matrix, and the delta matrix are denoted by $\mathbf{W}$, $\mathbf{W}_{0}$, and $\Delta\mathbf{W}$, respectively ($\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$).
+The activation vector of the previous layer is denoted by $\mathbf{x}\in\mathbb{R}^{d_{2}}$.
+The post-transformation vector is $\mathbf{z}$, where $\mathbf{z}=\mathbf{W}\mathbf{x}\in\mathbb{R}^{d_{1}}$, and the incremental part is denoted by $\Delta\mathbf{z}$, where $\Delta\mathbf{z}=\Delta\mathbf{W}\mathbf{x}\in\mathbb{R}^{d_{1}}$.
 The matrices $\mathbf{A}$ and $\mathbf{B}$ are low-rank matrices introduced by LoRA to represent $\Delta\mathbf{W}$, with $r$ being their rank.
-$r_{v}$ is used to specify the random projection matrix rank used in VeRA.
-The circular convolution kernel of C3A is denoted by $\Delta\mathbf{w}$.
+$r_{v}$ specifies the rank of the random projection matrix used in VeRA.
+The circular convolution kernel of C3A is denoted by $\Delta\mathbf{w}$ and the circular convolution operator by $\star$.
 The loss function is represented by $\mathcal{L}$.
@@
 The Hadamard product is denoted by $\circ$.
-In terms of dimensions, we have $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d\times d}$, and $\mathbf{x},\mathbf{z},\Delta\mathbf{z},\Delta\mathbf{w}\in\mathbb{R}^{d}$.
-The matrices $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d}$ are controlled by the rank $r$.
 3.2 Circular Convolution
-The circular convolution operator is defined as $\Delta\mathbf{z}=\Delta\mathbf{w}\star\mathbf{x}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, where $\mathcal{C}(\Delta\mathbf{w})$ is the circulant matrix of $\Delta\mathbf{w}$.
-The first row of $\mathcal{C}(\Delta\mathbf{w})$ is $\Delta\mathbf{w}$ and each row equals the row above it periodically shifted to right by one element.
-Concretely,
-$\mathcal{C}(\Delta\mathbf{w})=\begin{bmatrix}\Delta w_{1}&\Delta w_{2}&\cdots&\Delta w_{d-1}&\Delta w_{d}\\ \Delta w_{d}&\Delta w_{1}&\cdots&\Delta w_{d-2}&\Delta w_{d-1}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ \Delta w_{3}&\Delta w_{4}&\cdots&\Delta w_{1}&\Delta w_{2}\\ \Delta w_{2}&\Delta w_{3}&\cdots&\Delta w_{d}&\Delta w_{1}\\ \end{bmatrix}\in\mathbb{R}^{d\times d}.$
-Theoretically, the rank of $\mathcal{C}(\Delta\mathbf{w})$ is $d-\mathrm{Deg}(\mathrm{gcd}(f(x),x^{d}-1))$ Ingleton (1956), where $\mathrm{Deg}(\cdot)$ represents the degree of a polynomial, $f(x)$ is the associated polynomial of $\Delta\mathbf{w}$ (i.e., $f(x)=\sum_{i=1}^{d}\Delta w_{i}x^{i-1}$), and $\mathrm{gcd}(\cdot)$ means the greatest common divisor.
-By learning $\Delta\mathbf{w}$ in the $\mathbb{R}^{n}$ oracle, C3A automatically achieves dynamic rank selection that is not linearly bounded by the number of learnable parameters as LoRA.
-To achieve efficiency, we leverage the beautiful circulant structure of $\mathcal{C}(\Delta\mathbf{w})$, which makes it diagonalizable by the Fourier basis ($\mathbf{F}$) .
+Firstly, for simplicity, we assume $d_{1}=d_{2}=d$ and $\Delta\mathbf{w}\in\mathbb{R}^{d}$.
+The circular convolution operator is defined as $\Delta\mathbf{z}=\Delta\mathbf{w}\star\mathbf{x}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, where $\mathcal{C}(\cdot)$ is a function which takes a vector and outputs the corresponding circulant matrix.
+Concretely, the first row of $\mathcal{C}(\Delta\mathbf{w})$ is $\Delta\mathbf{w}$ and the following rows are equal to the row above them periodically shifted to the right by one element.
+In math,
+$\mathcal{C}(\Delta\mathbf{w})=\begin{bmatrix}\Delta w_{1}&\Delta w_{2}&\cdots&\Delta w_{d-1}&\Delta w_{d}\\ \Delta w_{d}&\Delta w_{1}&\cdots&\Delta w_{d-2}&\Delta w_{d-1}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ \Delta w_{3}&\Delta w_{4}&\cdots&\Delta w_{1}&\Delta w_{2}\\ \Delta w_{2}&\Delta w_{3}&\cdots&\Delta w_{d}&\Delta w_{1}\\ \end{bmatrix}.$
+Theoretically, the rank of $\mathcal{C}(\Delta\mathbf{w})$ is given by $d-\mathrm{Deg}(\mathrm{gcd}(f(x),x^{d}-1))$ Ingleton (1956), where $\mathrm{Deg}(\cdot)$ denotes the degree of a polynomial, $f(x)$ is the polynomial associated with $\Delta\mathbf{w}$ (i.e., $f(x)=\sum_{i=1}^{d}\Delta w_{i}x^{i-1}$), and $\mathrm{gcd}(\cdot)$ represents the greatest common divisor.
+Consequently, the theoretical upper bound on the rank of $\mathcal{C}(\Delta\mathbf{w})$ is $d$.
+By learning $\Delta\mathbf{w}$ in the $\mathbb{R}^{n}$ oracle, C3A automatically achieves dynamic rank selection, which is not linearly constrained by the number of learnable parameters, unlike LoRA.
+To achieve high efficiency, we leverage the beautiful circulant structure of $\mathcal{C}(\Delta\mathbf{w})$, which makes it diagonalizable by the Fourier basis ($\mathbf{F}$) .
 In math, it can be described as $\mathcal{C}(\Delta\mathbf{w})=\mathbf{F}\frac{\Lambda}{d}\mathbf{F}^{-1}$ Golub & Van Loan (1996), where $\Lambda$ is its eigenvalues and can be calculated by a Fourier transform of the first row (i.e., $\Lambda=\mathrm{diag}(\mathbf{F}\Delta\mathbf{w})$).
@@
 Since $\Delta\mathbf{z}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, it follows that $\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})$.
-For $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}$, we note that the circular convolution operation is commutative (i.e., $\Delta\mathbf{w}\star\mathbf{x}=\mathbf{x}\star\Delta\mathbf{w}$), thus $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})$.
+For $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}$, we note that the circular convolution operation is commutative (i.e., $\mathcal{C}(\Delta\mathbf{w})\mathbf{x}=\mathcal{C}(\mathbf{x})\Delta\mathbf{w}$), thus $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})$.
 Substituting these results into Equation 2, we obtain:
@@
 3.4 Block-Circular Convolution
-Despite the elegance and efficiency of the circular convolution operator, it is worth noting that we can not alter the number of learnable parameters due to the requirement of the convolution kernel being the same size as the activation vector.
-To enable this flexibility in case of difficult downstream tasks, we resort to the block-circular convolution.
-We can simply split the activation vector $\mathbf{x}$ (and the post-transformation vector $\mathbf{z}$) into blocks, and assign each pair of blocks with different convolution kernels.
+Despite the elegance and efficiency of the circular convolution operator, it has 2 fundamental limitations due to the requirement that the convolution kernel must be the same size as the activation vector: ❶ It cannot be applied to non-square weight matrices.The number of learnable parameters cannot be altered.
+The first limitation restricts its application in scenarios such as fine-tuning a LLaMA3-8B model, where the weight matrix sizes include $4096\times 1024$.
+The second limitation reduces the flexibility of C3A, making it challenging to handle complex downstream tasks that require a larger number of learnable parameters.
+To address these limitations, we resort to block-circular convolution.
+By splitting the activation vector $\mathbf{x}$ and the post-transformation vector $\mathbf{z}$ into blocks of the same size, we can assign unique convolution kernels to each pair of these blocks.
 To be concrete,
-$\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\cdots&\mathbf{x}_{b}\end{bmatrix}\qquad\Delta\mathbf{z}=\begin{bmatrix}\Delta\mathbf{z}_{1}&\Delta\mathbf{z}_{2}&\cdots&\Delta\mathbf{z}_{b}\end{bmatrix},$
-where $b$ is the block number. We will need $b^{2}$ convolution kernels to densely connect these blocks, which can be expressed as
-$\Delta\mathbf{z}_{i}=\sum_{j=1}^{b}\Delta\mathbf{w}_{ij}\star\mathbf{x}_{j},i\in\{1,2,\cdots,b\},$
-and can also be represented by a block-circular matrix
-$\Delta\mathbf{z}=\mathcal{C}_{b}(\Delta\mathbf{w})\mathbf{x}\qquad\mathcal{C}_{b}(\Delta\mathbf{w})=\begin{bmatrix}\mathcal{C}(\Delta\mathbf{w}_{11})&\mathcal{C}(\Delta\mathbf{w}_{12})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{1b})\\ \mathcal{C}(\Delta\mathbf{w}_{21})&\mathcal{C}(\Delta\mathbf{w}_{22})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{2b})\\ \cdots&\cdots&\cdots&\cdots\\ \mathcal{C}(\Delta\mathbf{w}_{b1})&\mathcal{C}(\Delta\mathbf{w}_{b2})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{bb})\end{bmatrix}.$
-In this case, $\Delta\mathbf{w}_{ij}\in\mathbb{R}^{\frac{d}{b}}$ (it is required that $d$ is divisible by $b$) and we have $b^{2}\frac{d}{b}=bd$ learnable parameters.
-$b$ acts as a hyperparameter controlling the number of learnable parameters here, just like $r$ for LoRA.
+$\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\cdots&\mathbf{x}_{\frac{d_{2}}{b}}\end{bmatrix},\qquad\Delta\mathbf{z}=\begin{bmatrix}\Delta\mathbf{z}_{1}&\Delta\mathbf{z}_{2}&\cdots&\Delta\mathbf{z}_{\frac{d_{1}}{b}}\end{bmatrix},$
+where $b$ is the block size and $b$ need to be a common divisor of $d_{1}$ and $d_{2}$.
+We will need $\frac{d_{1}d_{2}}{b^{2}}$ convolution kernels to densely connect these blocks, which can be expressed in math as
+$\Delta\mathbf{z}_{i}=\sum_{j=1}^{\frac{d_{2}}{b}}\Delta\mathbf{w}_{ij}\star\mathbf{x}_{j},i\in\{1,2,\cdots,\frac{d_{1}}{b}\}.$
+This calculation can be represented by a block-circular matrix:
+$\Delta\mathbf{z}=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{x},\qquad\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})=\begin{bmatrix}\mathcal{C}(\Delta\mathbf{w}_{11})&\mathcal{C}(\Delta\mathbf{w}_{12})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{1\frac{d_{2}}{b}})\\ \mathcal{C}(\Delta\mathbf{w}_{21})&\mathcal{C}(\Delta\mathbf{w}_{22})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{2\frac{d_{2}}{b}})\\ \cdots&\cdots&\cdots&\cdots\\ \mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}1})&\mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}2})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}\frac{d_{2}}{b}})\end{bmatrix}.$
+In this case, $\Delta\mathbf{w}_{ij}\in\mathbb{R}^{b}$ and we have $\frac{d_{1}d_{2}}{b^{2}}b=\frac{d_{1}d_{2}}{b}$ learnable parameters.
+In fact, $b$ acts as a hyperparameter controlling the number of learnable parameters, just like $r$ for LoRA.
 However, we highlight that while $r$ controls the rank of the delta matrix and learnable parameter number at the same time, $b$ only affects the latter.
@@
 3.5 Complexity Analysis
-We compare time compelxity and space complexity of LoRA, VeRA and C3A in Table 1. Detailed analysis follows.
+We compare the time complexity and space complexity of LoRA, VeRA and C3A in Table 1. Detailed analysis follows in this section.
+3.5.1 Time Complexity
+LoRA incorporates low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, which are sequentially multiplied with the activation vector, resulting in a time complexity of $\mathcal{O}(r(d_{1}+d_{2}))$.
+Typically, $r\ll\min(d_{1},d_{2})$.
+Conversely, VeRA, despite its high-rank nature and minimal number of trainable parameters, incurs a prohibitive time complexity of $\mathcal{O}(r_{v}(d_{1}+d_{2}))$, where $r_{v}$ can exceed $\max(d_{1},d_{2})$.
+Therefore, striking a balance between high rank and computational efficiency is never a low-hanging fruit.
+Thanks to the $\mathcal{O}(n\log n)$ complexity of the FFT algorithm used in Equation 1, C3A achieves a time complexity of $\mathcal{O}((d_{1}+d_{2})\log{b}+\frac{d_{1}d_{2}}{b})$.
+The first term is the time complexity for FFT, and the second term is for aggregation.
+In practical scenarios, $b$ is chosen as the greatest common divisor of $d_{1}$ and $d_{2}$ to achieve a high compression ratio.
+Given that, C3A is comparable to LoRA in terms of time complexity.
 Table 1: Time and space complexity comparison of LoRA, VeRA and C3A.
 We split the space complexity into Parameter number and Other auxiliary tensors to help better understand the differences.
-We highlight that in practice, to achieve similar performance, $b\leq r\ll r_{v}$.
-Method Time # Parameter # Other # Total
-LoRA $\mathcal{O}(rd)$ $2rd$ $r$ $2rd+r$
-VeRA $\mathcal{O}(r_{v}d)$ $r_{v}+d$ $2r_{v}d+r_{v}+d$ $2(r_{v}d+r_{v}+d)$
-C3A $\mathcal{O}(bd\log\frac{d}{b})$ $bd$ 00 $bd$
-3.5.1 Time Complexity
-LoRA incorporates low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, which are sequentially multiplied with the activation vector, resulting in a time complexity of $\mathcal{O}(rd)$.
-Typically, $r\ll d$.
-Due to the $\mathcal{O}(d\log d)$ complexity of the FFT algorithm used in Equation 1, C3A achieves a time complexity of $\mathcal{O}(bd\log\frac{d}{b})$.
-Given that $b\ll d$, C3A is comparable to LoRA in terms of time complexity.
-Conversely, VeRA, despite its high-rank nature and minimal number of trainable parameters, incurs a prohibitive time complexity of $\mathcal{O}(r_{v}d)$, where $r_{v}$ can exceed $d$.
-Therefore, striking a balance between high rank and computational efficiency is never a low-hanging fruit.
+We highlight that in practice, to achieve similar performance, $\frac{\max(d_{1},d_{2})}{b}\leq r\ll r_{v}$.
+Method Time # Param # Other # Total
+LoRA $\mathcal{O}(r(d_{1}+d_{2}))$ $r(d_{1}+d_{2})$ 0 $r(d_{1}+d_{2})$
+VeRA $\mathcal{O}(r_{v}(d_{1}+d_{2}))$ $r_{v}+d_{1}$ $r_{v}(d_{1}+d_{2})$ $r_{v}(d_{1}+d_{2})+r_{v}+d_{1}$
+C3A $\mathcal{O}((d_{1}+d_{2})\log{b}+\frac{d_{1}d_{2}}{b})$ $\frac{d_{1}d_{2}}{b}$ $b$ $\frac{d_{1}d_{2}}{b}+b$
 3.5.2 Space Complexity
 We analyze the space complexity of LoRA, VeRA, and C3A during training.
-The differences among these methods primarily arise from the trainable parameter size and the auxiliary tensors required for the forward pass and backpropagation.
-For LoRA, an intermediate tensor of size $r$ is introduced (i.e., $\mathbf{Ax}\in\mathbb{R}^{r}$).
-In contrast, VeRA necessitates two random projection matrices, with a total size of $2r_{v}d$.
-Additionally, VeRA introduces two intermediate tensors of sizes $d$ and $r_{v}$.
-Thanks to the efficient backpropagation algorithm detailed in Section 3.3, C3A does not require any additional intermediate tensors.
-This optimization significantly reduces its space complexity compared to LoRA and VeRA.
+The differences among these methods primarily arise from the trainable parameters and the auxiliary tensors required for the forward pass and backpropagation.
+LoRA does not rely on auxiliary tensors, while VeRA necessitates 2 random projection matrices, with a total size of $r_{v}(d_{1}+d_{2})$.
+Since $r_{v}$ is by no means negligible, the memory consumption of VeRA would be significantly larger than that of LoRA.
+In terms of C3A, the only additional auxiliary tensor would be of size $b\leq\min(d_{1},d_{2})$, which is reserved by the FFT algorithm.
+Given a proper $b$, which is often close to the greatest common divisor of $d_{1}$ and $d_{2}$, the space complexity of C3A excels.
+Additionally, we observe that when applying FFT, input blocks and output blocks are independent within their own group.
+Therefore, parallelization can be employed to enable the flexibility of trading space for time.
 4 Experiment
-In this section, we comprehensively evaluate C3A in both natural language processing (NLP) and computer vision (CV).
+In this section, we firstly conduct experiments on a synthetic dataset to show the superior expressiveness of C3A over LoRA.
+After that, we comprehensively evaluate C3A in both natural language processing (NLP) and computer vision (CV).
 For NLP, we first demonstrate the effectiveness of C3A using the GLUE benchmark with RoBERTa-Base and RoBERTa-Large models.
 Additionally, we apply C3A to fine-tuning the LLaMA family models, which stands as a representative of prevalent large language models (LLMs).
-For CV, we examine two practical scenarios: the traditional classification task using Vision Transformers (ViTs) and the more intricate text-to-image generation task based on diffusion models.
+For CV, we examine the most representative classification task using Vision Transformers (ViTs) on a wide range of datasets.
+4.1 Synthetic Data
+Settings.
+We distribute 8 points evenly on a 2D plane as cluster centers and randomly sample 30 points from 8 corresponding multivariate Gaussian distributions.
+A 3-layer MLP is then used to classify these point clusters.
+To compare the expressiveness of 2 types of layers, we replace the middle layer with either a low-rank layer or a circulant layer, ensuring that both layers have the same number of trainable parameters for a fair comparison.
+Figure 3: Expressiveness test on synthetic data.
+The left figure shows the synthetic data used for the experiment, while the right figure illustrates the training accuracy curves of a 3-layer MLP, incorporating C3A, LoRA, and standard linear layers, respectively.
+Results.
+The results are presented in Figure 3.
+We observe that LoRAr=1 struggles with this straightforward classification task.
+In contrast, C3Ab=128/2, despite using the same number of parameters, achieves perfect classification, comparable to a standard linear layer.
+This demonstrates the high expressiveness of C3A given the same parameter budget.
 4.2 Natural Language Understanding
@@
 LoRA is a widely known PEFT method that employs low-rank decomposition to compress additive delta matrices.
-VeRA is a recent approach that focuses on further reducing trainable parameters for LoRA.
+VeRA is a recent approach that focuses on further reducing trainable parameters of LoRA while preserving a high rank.
 BOFT is another innovative method in PEFT research, compressing multiplicative delta matrices using orthogonal decomposition and butterfly factorization.
+Table 2: Performance of different PEFT methods on the GLUE benchmark.
+We fine-tune pre-trained RoBERTa-Base and -Large models on 6 datasets.
+We report the Matthew’s Correlation Coefficient (MCC) for CoLA, Pearson Correlation Coefficient (PCC) for STS-B, and accuracy (Acc.) for all the remaining tasks.
+For each metric, a higher score indicates better performance.
+“Avg.” denotes the average score of each method across all datasets.
+The best results for each dataset are highlighted in bold.
+# Trainable parameters does not include the classification head since each method uses a head of the same size.
+Memory Cost is measured on fixed length (i.e., $256$) data with a batchsize of $64$.
+Method # Trainable Parameters Memory Cost (GB) SST-2 MRPC CoLA QNLI RTE STS-B Avg.
+Base Full 124M 17.19 94.01$\pm$0.39 87.10$\pm$0.79 62.00$\pm$1.16 92.40$\pm$0.28 77.33$\pm$2.68 90.70$\pm$0.14 83.92
+BitFit 102K 12.60 93.30$\pm$0.30 85.80$\pm$0.21 59.21$\pm$1.74 91.96$\pm$0.18 73.07$\pm$1.34 90.18$\pm$0.17 82.25
+(IA)3  [^1] K 19.86 92.98$\pm$0.34 85.86$\pm$0.59 60.49$\pm$1.09 91.56$\pm$0.17 69.10$\pm$1.18 90.06$\pm$0.21 81.67
+LoRAr=8 295K 13.75 94.50$\pm$0.41 85.68$\pm$0.74 60.95$\pm$1.57 92.54$\pm$0.20 76.68$\pm$1.42 89.76$\pm$0.39 83.35
+VeRAr=1024 43K 15.51 93.97$\pm$0.17 86.23$\pm$0.41 62.24$\pm$1.91 91.85$\pm$0.17 75.74$\pm$1.56 90.27$\pm$0.25 83.38
+[…]
+C3Ab=768/1 18K 12.83 93.42$\pm$0.26 86.33$\pm$0.32 61.83$\pm$0.96 91.83$\pm$0.04 76.17$\pm$1.39 90.46$\pm$0.29 83.34
+C3Ab=768/6  [^1] K 12.72 94.20$\pm$0.16 86.67$\pm$0.54 62.48$\pm$1.20 92.32$\pm$0.25 77.18$\pm$1.41 90.16$\pm$0.42 83.84
+Large BitFit 271K 30.65 95.09$\pm$0.27 88.10$\pm$0.76 65.40$\pm$0.76 94.06$\pm$0.14 82.60$\pm$1.15 91.73$\pm$0.20 86.16
+(IA)3 295K 48.81 95.32$\pm$0.20 87.06$\pm$0.57 66.52$\pm$1.10 94.18$\pm$0.15 84.33$\pm$2.38 91.58$\pm$0.39 86.50
+LoRAr=8 786K 34.12 95.53$\pm$0.35 86.12$\pm$3.22 65.16$\pm$0.76 93.73$\pm$0.30 83.75$\pm$0.51 91.46$\pm$0.21 85.96
+VeRAr=256 61K 34.16 95.83$\pm$0.43 87.72$\pm$0.55 63.66$\pm$1.45 94.11$\pm$0.20 83.03$\pm$1.65 91.12$\pm$0.37 85.91
+[…]
+C3Ab=1024/1 49K 31.83 95.78$\pm$0.05 88.02$\pm$0.62 66.59$\pm$1.20 94.22$\pm$0.25 82.89$\pm$0.67 91.86$\pm$0.14 86.56
+C3Ab=1024/8 393K 31.79 95.78$\pm$0.15 88.09$\pm$0.47 67.18$\pm$1.92 94.26$\pm$0.19 84.62$\pm$1.36 91.81$\pm$0.36 86.96
 Settings.
@@
 For the unique hyperparameters of different baselines, we adopt the values suggested in the original papers (e.g., VeRA’s $r$ and BOFT’s $b$ and $m$).
-The number of trainable parameters excludes the classification head, as each method uses the same one.
-The shared hyperparameters (i.e., the learning rate for classification head and the learning rate for other trainable parameters) are found by hyperparameter search.
+The number of trainable parameters excludes the classification head, as each method uses one of the same size.
+The shared hyperparameters (i.e., the learning rate for classification head and for other trainable parameters, separately) are found by hyperparameter search.
 For the memory cost, to ensure fairness and consistency, we fix the length of input data to 256 tokens and use a batchsize of 64.
@@
 Results are summarized in Table 2.
-In general, C3Ab=1 achieves better or on-par performance compared to baseline methods with extremely small trainable parameters.
-When a larger number of trainable parameters is allowed, C3Ab=6 and C3Ab=8 significantly outperform the baselines.
-Furthermore, compared to (IA)3, LoRA, VeRA, and BOFT, despite training a similar number of or more parameters, C3A stands out for its excellent memory efficiency.
+In general, C3Ab=768/1 and C3Ab=1024/1 achieves better or on-par performance compared to baseline methods with extremely small trainable parameters.
+When a larger number of trainable parameters is allowed, C3Ab=768/6 and C3Ab=1024/8 significantly outperform baselines.
+In addition, compared to (IA)3, LoRA, VeRA, and BOFT, despite training a similar number of or more parameters, C3A stands out for its excellent memory efficiency.
 The only method with better memory efficiency is BitFit, which can be seen as an upper bound since it does not introduce any new parameters.
-Table 2: Performance of different PEFT methods on the GLUE benchmark.
-We fine-tune pre-trained RoBERTa-Base and -Large models on 6 datasets.
-We report the Matthew’s correlation coefficient (MCC) for CoLA, Pearson correlation coefficient (PCC) for STS-B, and accuracy (Acc.) for all remaining tasks.
-For all 6 datasets, a higher metric indicates better performance.
+Furthermore, nearly all circulant delta matrices discovered by C3A are full rank, indicating maximal capacity Zeng & Lee (2023) and providing theoretical support for the impressive results.
+4.3 Instruction Tuning
+Settings.
+We evaluate C3A against LoRA by fine-tuning LLaMA2-7B/13B and LLaMA3-8B on 4 standard instruction tuning datasets: GSM8k Cobbe et al.
+(2021), MATH Yu et al.
+(2023), ViGGO Juraska et al.
+(2019), and SQL Zhong et al.
+(2017), following the setting outlined in Nikdan et al.
+(2024).
+For the SQL dataset, we preprocess by selecting 12.5% of the dataset, followed by an 4:1 train-test split.
+To ensure a fair comparison, we maintain LoRA parameters with $r=32$ and a dropout rate of 0.05, while exploring various $\alpha$ values ranging from 32 to 128.
+All experiments are conducted for a standard single pass (epoch) over the dataset.
+Results.
+In Table 3, we summarize our main experiments.
+It can be observed that C3A generally surpasses LoRA across both LLaMA2-7B/13B and LLaMA3-8B.
+Notably, C3A only utilizes less than half of the parameter count compared to LoRA.
+Overall, these results clearly highlight the effectiveness of C3A.
+Table 3: Comparison of C3A and LoRA on fine-tuning LLaMA2 and LLaMA3 models in terms of accuracy and trainable parameters.
+The best results for each dataset are highlighted in bold.
 “Avg.” denotes the average accuracy of each method across all datasets.
+Model Method # Trainable Parameters GSM8k MATH ViGGO SQL Avg.
+LLaMA2-7B LoRAr=32 16.8M 27.22 4.10 92.24 65.39 47.24
+C3Ab=4096/32 8.4M 29.34 4.25 94.03 66.21 48.46
+LLaMA2-13B LoRAr=32 26.2M 42.68 6.70 93.19 67.53 52.53
+C3Ab=5120/32 13.1M 42.00 6.65 95.57 67.28 52.88
+LLaMA3-8B LoRAr=32 13.6M 56.57 18.95 91.78 68.19 58.87
+C3Ab=4096/32 5.2M 57.54 19.75 93.07 69.47 59.96
+4.4 Image Classification
+Settings.
+We focus on the task of image classification utilizing the Vision Transformer (ViT) models.
+Specifically, we employ both the Base and Large variants of this prevalent computer vision foundation model, as introduced by Dosovitskiy et al.
+(2020).
+These ViT models are pre-trained on the extensive ImageNet-21K dataset Ridnik et al.
+(2021).
+For the fine-tuning phase, we utilize a diverse set of datasets including Pets, Cars, DTD, EuroSAT, FGVC and RESISC.
+Table 4: Fine-tuning results with ViT-Base and ViT-Large models on various image classification datasets.
+The models are fine-tuned for 10 epochs, and the best-performing model, based on validation set accuracy, is selected.
+The reported accuracy corresponds to the performance on the test set.
 The best results for each dataset are highlighted in bold.
-# Trainable parameters does not include the classification head since each method uses the same head.
-Memory Cost is measured on fixed length (i.e., $256$) data with a batchsize of $64$.
-Methods # Trainable Parameters Memory Cost (GB) SST-2 MRPC CoLA QNLI RTE STS-B Avg.
-Base BitFit 102K 12.60 93.30$\pm$0.30 85.80$\pm$0.21 59.21$\pm$1.74 91.96$\pm$0.18 73.07$\pm$1.34 90.18$\pm$0.17 82.25
-(IA)3  [^1] K 19.86 92.98$\pm$0.34 85.86$\pm$0.59 60.49$\pm$1.09 91.56$\pm$0.17 69.10$\pm$1.18 90.06$\pm$0.21 81.67
-LoRAr=8 295K 13.75 94.50$\pm$0.41 85.68$\pm$0.74 60.95$\pm$1.57 92.54$\pm$0.20 76.68$\pm$1.42 89.76$\pm$0.39 83.35
-VeRAr=1024 43K 15.51 93.97$\pm$0.17 86.23$\pm$0.41 62.24$\pm$1.91 91.85$\pm$0.17 75.74$\pm$1.56 90.27$\pm$0.25 83.38
-[…]
-C3Ab=1 18K 12.83 93.42$\pm$0.26 86.33$\pm$0.32 61.83$\pm$0.96 91.83$\pm$0.04 76.17$\pm$1.39 90.46$\pm$0.29 83.34
-C3Ab=6  [^1] K 12.72 94.20$\pm$0.16 86.67$\pm$0.54 62.48$\pm$1.20 92.32$\pm$0.25 77.18$\pm$1.41 90.16$\pm$0.42 83.84
-Large BitFit 271K 30.65 95.09$\pm$0.27 88.10$\pm$0.76 65.40$\pm$0.76 94.06$\pm$0.14 82.60$\pm$1.15 91.73$\pm$0.20 86.16
-(IA)3 295K 48.81 95.32$\pm$0.20 87.06$\pm$0.57 66.52$\pm$1.10 94.18$\pm$0.15 84.33$\pm$2.38 91.58$\pm$0.39 86.50
-LoRAr=8 786K 34.12 95.53$\pm$0.35 86.12$\pm$3.22 65.16$\pm$0.76 93.73$\pm$0.30 83.75$\pm$0.51 91.46$\pm$0.21 85.96
-VeRAr=256 61K 34.16 95.83$\pm$0.43 87.72$\pm$0.55 63.66$\pm$1.45 94.11$\pm$0.20 83.03$\pm$1.65 91.12$\pm$0.37 85.91
-[…]
-C3Ab=1 49K 31.83 95.78$\pm$0.05 88.02$\pm$0.62 66.59$\pm$1.20 94.22$\pm$0.25 82.89$\pm$0.67 91.86$\pm$0.14 86.56
-C3Ab=8 393K 31.79 95.78$\pm$0.15 88.09$\pm$0.47 67.18$\pm$1.92 94.26$\pm$0.19 84.62$\pm$1.36 91.81$\pm$0.36 86.96
+“Avg.” denotes the average accuracy of each method across all datasets.
+Method # Trainable Parameters Pets Cars DTD EuroSAT FGVC RESISC Avg.
+Base Head - 90.28$\pm$0.43 25.76$\pm$0.28 69.77$\pm$0.67 88.72$\pm$0.13 17.44$\pm$0.43 74.22$\pm$0.10 61.03
+Full 85.8M 92.82$\pm$0.54 85.10$\pm$0.21 80.11$\pm$0.56 99.11$\pm$0.07 61.60$\pm$1.00 96.00$\pm$0.23 85.79
+LoRAr=16 590K 93.76$\pm$0.44 78.04$\pm$0.33 78.56$\pm$0.62 98.84$\pm$0.08 56.64$\pm$0.55 94.66$\pm$0.17 83.42
+C3Ab=768/12 221K 93.88$\pm$0.22 79.05$\pm$0.35 80.57$\pm$0.53 98.88$\pm$0.07 54.31$\pm$0.79 94.54$\pm$0.23 83.54
+Large Head - 91.11$\pm$0.30 37.91$\pm$0.27 73.33$\pm$0.26 92.64$\pm$0.08 24.62$\pm$0.24 82.02$\pm$0.11 66.94
+Full 303M 94.30$\pm$0.31 88.15$\pm$0.50 80.18$\pm$0.66 99.06$\pm$0.10 67.38$\pm$1.06 96.08$\pm$0.20 87.53
+LoRAr=16 1.57M 94.62$\pm$0.47 86.11$\pm$0.42 80.09$\pm$0.42 98.99$\pm$0.03 63.64$\pm$0.83 95.94$\pm$0.21 86.56
+C3Ab=1024/16 786K 94.48$\pm$0.30 84.94$\pm$0.39 82.62$\pm$0.52 98.75$\pm$0.19 63.80$\pm$0.37 95.52$\pm$0.16 86.69
+Results.
+Table 4 presents a summary of the results obtained from 6 image classification datasets using the ViT Base and Large models.
+The LoRA and C3A methods both show substantial performance improvements over Head Tuning, highlighting their effectiveness in the image classification domain.
+Notably, our method achieves comparable performance to LoRA while utilizing only half of the parameter count.
 5 Conclusion
@@
 Consequently, C3A presents itself as a compelling alternative to LoRA for model fine-tuning.
-Appendix A Appendix

### v3

@@
 Extensive experiments demonstrate that C3A consistently outperforms LoRA and its variants across various fine-tuning tasks.
+Code released.
+Parameter-Efficient Fine-Tuning via Circular Convolution
+Aochuan Chen[Uncaptioned image]thanks: Equal contribution.
+Jiashun Cheng[Uncaptioned image]11footnotemark: 1 Zijing Liu[Uncaptioned image] Ziqi Gao[Uncaptioned image] Fugee Tsung[Uncaptioned image] Yu Li[Uncaptioned image] Jia Li[Uncaptioned image]thanks: Corresponding author.
+[Uncaptioned image]The Hong Kong University of Science and Technology (Guangzhou) The Hong Kong University of Science and Technology [Uncaptioned image]International Digital Economy Academy jialee@hkust-gz.edu.cn
 1 Introduction
-Figure 1: Relative comparison of C3A and baselines on RoBERTa-Base.
-The Pearson Correlation Coefficient (PCC) is evaluated on STS-B and the Matthew’s Correlation Coefficient (MCC) on CoLA.
-Accuracies across SST-2, MRPC, QNLI, and RTE are averaged and reported as Acc.
-#Param denotes the count of learnable parameters (excluding the head), while Memory indicates the peak GPU memory usage during training.
-Metrics highlighted in blue represent values where higher is better, and those in red indicate values where lower is better.
-Refer to Table 2 for absolute statistics.
-In recent years, Large Foundation Models (LFMs) have surged in popularity due to their remarkable performance across various tasks in natural language processing (NLP) Brown et al.
-(2020); Touvron et al.
-(2023), computer vision (CV) Radford et al.
-(2021); Kirillov et al.
-(2023), and other domains Li et al.
-(2024).
-Characterized by their vast number of parameters and substantial computational demands, these models have set new benchmarks in accuracy and versatility.
-However, their size and complexity pose significant challenges for efficient fine-tuning, particularly in resource-constrained environments Malladi et al.
-(2023); Zhang et al.
-(2024b).
-To address this issue, Parameter-efficient fine-tuning (PEFT) techniques Mangrulkar et al.
-(2022), such as Low-Rank Adaptation (LoRA) Hu et al.
-(2021), have emerged as outstanding practical solutions.
-LoRA reduces the number of trainable parameters by employing low-rank matrices to approximate weight changes, making fine-tuning more feasible and accessible without compromising the model’s performance.
-Concretely, LoRA can be expressed in math:
+Figure 1: Performance comparison of C3A and other methods relative to LoRA on LLaMA-8B. Higher values are better across all metrics. See Table 3 and Table 4 for more statistics.
+In recent years, Large Foundation Models (LFMs) have witnessed a pronounced ascendance in both scholarly and practical realms, attributable to their exceptional efficacy across diverse tasks in natural language processing (NLP) (Brown et al., 2020; Touvron et al., 2023), computer vision (CV) (Radford et al., 2021; Kirillov et al., 2023), and other domains (Li et al., 2024a, b, 2025).
+Distinguished by an extensive parameter count and significant computational requisites, these models have established unprecedented benchmarks in both accuracy and versatility.
+Nonetheless, their considerable size and intricate structure present formidable obstacles for efficient fine-tuning, especially within resource-constrained environments (Malladi et al., 2023; Zhang et al., 2024b).
+To mitigate these challenges, parameter-efficient fine-tuning (PEFT) techniques (Mangrulkar et al., 2022), exemplified by Low-Rank Adaptation (LoRA) (Hu et al., 2021), have emerged as highly effective solutions.
+LoRA reduces the number of trainable parameters by leveraging low-rank matrices to approximate alterations in weights, thereby facilitating fine-tuning without degrading the model’s efficacy.
+Specifically, LoRA can be articulated mathematically as follows:
 $\mathbf{Wx}=(\mathbf{W}_{0}+\Delta\mathbf{W})\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B(Ax)},$
-where $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ are weight matrices, $\mathbf{B}\in\mathbb{R}^{d_{1}\times r},\mathbf{A}\in\mathbb{R}^{r\times d_{2}}$ are low-rank matrices to construct $\Delta\mathbf{W}$ and $\mathbf{x}\in\mathbb{R}^{d_{2}}$ is the activations.
-The number of trainable parameters is $r(d_{1}+d_{2})$, prompting the choice $r\ll\min(d_{1},d_{2})$ (e.g., $r=8$ for $d_{1}=d_{2}=1024$) for achieving high parameter efficiency.
-However, as discussed in Zeng & Lee (2023), LoRA’s capacity to represent a target model is strictly limited by $r$.
-To address this trade-off between performance and efficiency, Kopiczko et al.
-(2023) proposed Vector Random Matrix Adaptation (VeRA), which achieves competitive performance with significantly fewer trainable parameters through fixed random-matrix projections.
-Despite its low parameter count, VeRA requires substantial computational power and memory due to the large size of the random matrices used for projection.
-This raises the following open research question in the context of PEFT:
-How can high-rank adaptation be achieved without incurring significant costs in terms of time complexity, memory usage, and trainable parameters?
-As discussed in Dosovitskiy et al.
-(2020), dense linear layers lack inductive biases, which poses challenge on the optimization process and therefore limits transformers’ ability to compete with Convolutional Neural Networks (CNNs) when data is scarce.
-Given the constraints of a small training set for the downstream task, we posit that incorporating a stronger inductive bias could improve the adaptation performance.
-In light of this, we propose Circular Convolution Adaptation (C3A), featuring the circular convolution operator Bamieh (2018).
+where $\mathbf{W},\mathbf{W}_{0},\Delta\mathbf{W}\in\mathbb{R}^{d_{1}\times d_{2}}$ are weight matrices, $\mathbf{B}\in\mathbb{R}^{d_{1}\times r},\mathbf{A}\in\mathbb{R}^{r\times d_{2}}$ are low-rank matrices formulated to construct $\Delta\mathbf{W}$, and $\mathbf{x}\in\mathbb{R}^{d_{2}}$ are the activations.
+The number of trainable parameters is $r(d_{1}+d_{2})$, thereby motivating the selection of $r\ll\min(d_{1},d_{2})$ (e.g., $r=8$ for $d_{1}=d_{2}=1024$) to attain elevated parameter efficiency.
+Nonetheless, as elaborated by Zeng and Lee (2023), the potential of LoRA to encapsulate a target model is inherently constrained by $r$.
+In an effort to reconcile the dichotomy between performance and efficiency, Kopiczko et al.
+(2023) introduced Vector Random Matrix Adaptation (VeRA).
+VeRA attains comparable performance with a markedly reduced count of trainable parameters via fixed random-matrix projections.
+However, despite its minimal parameter count, VeRA demands considerable computational resources and memory capacity due to the extensive nature of the random matrices employed for projection.
+As depicted in Figure 1, other representative works share the same resource problem.
+This precipitates the following open research question within the scope of PEFT:
+Beyond low parameter counts, how to achieve high-rank adaptation without incurring significant costs of time and memory?
+To address this question, we introduce Circular Convolution Adaptation (C3A), which incorporates the circular convolution operator (Bamieh, 2018).
+Circular convolution has garnered significant attention in both signal processing (Li et al., 2020) and cryptography (Dworkin et al., 2001) due to its exceptional efficiency and compactness.
+This operator can be equivalently expressed as multiplication by a circulant matrix, providing rank flexibility that is independent of the number of trainable parameters.
+Furthermore, by employing the Fast Fourier Transform (FFT), C3A achieves superior time and memory efficiency compared to the direct multiplication (Bamieh, 2018), which makes it competitive with LoRA in terms of efficiency.
+In addition, as explicated by Dosovitskiy et al.
+(2020), dense linear layers exhibit a deficiency of inductive biases, engendering a complex optimization landscape.
+Consequently, this hampers the effectiveness of transformers in comparison to Convolutional Neural Networks (CNNs) under conditions of limited data availability.
+Within the framework of a constrained training dataset for the downstream task, we postulate that a robust inductive bias could potentially augment adaption performance.
+The circular pattern in C3A serves precisely as such an inductive bias.
+In summary, circular convolution presents a promising solution for circumventing the rank limitations of LoRA at minimal costs. Our contributions can be summarized as follows:
+❶ We introduce C3A, a novel approach for PEFT.
+This method leverages the circular convolution operation and its equivalent circulant matrix to provide a flexible rank, which is free of linear constraint by the number of trainable parameters.
+❷ Leveraging the elegant diagonalization of the circulant matrix, we implement both the forward pass and backpropagation using FFT.
+With the incorporation of FFT, the computation and memory efficiency of C3A excels.
+C3A strikes a unique balance between performance an efficiency.
+❸ To offer greater flexibility in controlling the number of trainable parameters, we extend C3A by incorporating block-circular convolution, which results in block-circulant matrices.
+This extension allows C3A to achieve fully customizable parameter counts as well as adaptable rank configurations.
+❹ We validate C3A through comprehensive fine-tuning experiments across diverse tasks, which demonstrate C3A’s outstanding accuracy and memory merits compared to existing methods.
+2 Related Work
+2.1 Parameter-Efficient Fine-Tuning
+Research on PEFT has generally progressed along three main directions.
+The first direction involves partially updating the pre-trained neural network (e.g., the layer norm (Basu et al., 2024) or the biases (Zaken et al., 2021)).
+Traditional methods relied on hand-crafted heuristics (Raghu et al., 2019) to identify which parameters are crucial and should be fine-tuned.
+More advanced approaches employ optimization techniques (Guo et al., 2020; Xu et al., 2021; Fu et al., 2023).
+For example, Guo et al.
+(2020) reformulated such a discrete optimization problem into a continuous one by employing Bernoulli masks and the Gumbel-softmax approximation (Jang et al., 2016).
+The second direction emerged to maintain the integrity of the pre-trained model while enabling a high degree of parameter sharing through adapter-based methods (He et al., 2021; Rebuffi et al., 2017; Rücklé et al., 2020; Liu et al., 2022; Lian et al., 2022).
+These works focus on integrating additional modules, termed adapters, to fit the downstream task, effectively decoupling the pre-trained model parameters from those specific to the downstream task.
+Prompt Tuning (Brown et al., 2020; Gao et al., 2020; Chen et al., 2023; Zhang et al., 2024a) and Prefix Tuning (Li and Liang, 2021; Jia et al., 2022) also fall into this category, despite ignoring potential semantic meanings.
+The final direction is characterized by delta-weight-based methods, such as Low-Rank Adaptation (LoRA) (Hu et al., 2021) and Orthogonal Fine-tuning (OFT) (Qiu et al., 2023).
+These methods bridge the gap between the pre-trained model and the downstream task by adaptive delta weights, which are stored separately while used in combination with the pre-trained weights.
+This unique design enables disentanglement of the pretrained and downstream-specific weights.
+Namely, it achieves parameter sharing and preserves the ability to integrate without additional inference cost.
+LoRA models the delta-weights by an additive matrix while OFT does it by a multiplicative one.
+To further improve either parameter efficiency or performance, many variants has been proposed for both of the methods (Kopiczko et al., 2023; Liu et al., 2024b, 2023; Yuan et al., 2024; Hayou et al., 2024b).
+However, these methods can hardly achieve high parameter efficiency and performance without incurring heavy computation and memory usage.
+2.2 Circular Convolution
+Circular convolution has been extensively studied in signal processing (Rabiner et al., 1978; McGillem and Cooper, 1984; Li et al., 2020) and cryptography (Dworkin et al., 2001; Gong et al., 2024).
+Owing to its computational advantages, circular convolution has also been explored in machine learning for generating long embeddings of high-dimensional data (Yu et al., 2014) and compressing heavily parameterized layers (Cheng et al., 2015; Ding et al., 2017).
+Remarkably, it achieves these efficiencies without significant performance degradation, which makes it a promising technique for fine-tuning applications.
+Despite its success in small neural networks such as LeNet (Cheng et al., 2015), circular convolution has not demonstrated lossless performance in modern large foundational models (LFMs) or even in their base architecture, the transformer.
+This limitation may be attributed to the conflict between its high intrinsic bias (i.e., the circulant pattern) and the vast amount of data required for training LFMs.
+Conversely, when fine-tuning LFMs, it is often impractical to collect as much data as needed for training from scratch.
+In such scenarios, the intrinsic bias of circular convolution could potentially serve as a regularization mechanism, thereby benefiting the optimization process of fine-tuning.
+3 Method
 Figure 2: Overview of LoRA (A) and our C3A (B,C) method.
@@
 The total trainable parameter count is $\frac{d_{1}d_{2}}{b}$, which disentangles with the rank of the delta weight.
-Here, $b$ is the block size of the block-circular matrix and it should be a common divisor (CD) of $d_{1}$ and $d_{2}$.
-Circular convolution has been famous in signal processing Li et al.
-(2020) and cryptography Dworkin et al.
-(2001) for its distinguished efficiency and compactness.
-This operator can be equivalently represented as multiplication by a circulant matrix, which offers rank flexibility that is unconstrained by the number of trainable parameters.
-Moreover, leveraging the Fast Fourier Transform (FFT), C3A ensures better time and memory efficiency than directly multiplying the circulant matrix Bamieh (2018), rendering it competitive with LoRA from the efficiency aspect.
-In a word, circular convolution is a promising option for escaping the rank curse of LoRA at minimal cost.
-We summarize our contribution as follows:
-① We introduce C3A, a novel approach for PEFT using convolution.
-This method leverages the circular convolution operation and its equivalent circulant matrix to provide a flexible rank, which is free of linear constraint by the number of trainable parameters, for the delta matrix.
-② Leveraging the elegant diagonalization of the circulant matrix, we implement both the forward pass and backpropagation using FFT.
-With the incorporation of FFT, the computation and memory efficiency of C3A excels.
-C3A strikes a unique balance between performance an efficiency.
-③ To offer greater flexibility in controlling the number of trainable parameters, we extend C3A by incorporating block-circular convolution, which results in block-circulant matrices.
-This extension allows C3A to achieve fully customizable parameter counts as well as adaptable rank configurations.
-④ We validate C3A through comprehensive fine-tuning experiments across diverse tasks including natural language understanding, instruction tuning and image classification.
-Experiments demonstrate C3A’s outstanding accuracy and memory merits compared to existing state-of-the-art methods.
-2 Related Work
-2.1 Parameter-Efficient Fine-Tuning
-Research on Parameter-Efficient Fine-Tuning (PEFT) has generally progressed along three main directions.
-The first direction involves partially updating the pre-trained neural network (e.g., the body Oh et al.
-(2020) or the biases Zaken et al.
-(2021)).
-Traditional methods relied on hand-crafted heuristics Han et al.
-(2015); Oh et al.
-(2020) to identify which parameters are crucial and should be fine-tuned.
-More advanced approaches employ optimization techniques Guo et al.
-(2020); Xu et al.
-(2021); Fu et al.
-(2023).
-For example, Guo et al.
-(2020) reformulated such a discrete optimization problem into a continuous one by employing Bernoulli masks and the Gumbel-softmax approximation.
-The second direction emerged to maintain the integrity of the pre-trained model while enabling a high degree of parameter sharing through adapter-based methods He et al.
-(2021); Rebuffi et al.
-(2017); Rücklé et al.
-(2020); Liu et al.
-(2022).
-These works focus on integrating additional modules, termed adapters, to fit the downstream task, effectively decoupling the pre-trained model parameters from those specific to the downstream task.
-Prompt Tuning Brown et al.
-(2020); Gao et al.
-(2020); Chen et al.
-(2023); Zhang et al.
-(2024a) and Prefix Tuning Li & Liang (2021) also fall into this category, despite ignoring potential semantic meanings.
-The final direction is characterized by delta-weight-based methods, such as Low-Rank Adaptation (LoRA) Hu et al.
-(2021) and Orthogonal Fine-tuning (OFT) Qiu et al.
-(2023).
-These methods bridge the gap between the pre-trained model and the downstream task by adaptive delta weights, which are stored separately while used in combination with the pre-trained weights.
-This unique design enables disentanglement of the pretrained and downstream-specific weights.
-Namely, it achieves parameter sharing and preserves the ability to integrate without additional inference cost.
-LoRA models the delta-weights by an additive matrix while OFT does it by a multiplicative one.
-To further improve either parameter efficiency or performance, many variants has been proposed for both of the methods Kopiczko et al.
-(2023); Liu et al.
-(2024; 2023); Yuan et al.
-(2024).
-However, these methods can hardly achieve high parameter efficiency and performance without incurring heavy computation and memory usage.
-2.2 Circular Convolution
-Circular convolution has been extensively studied in signal processing Rabiner et al.
-(1978); McGillem & Cooper (1984); Li et al.
-(2020) and cryptography Dworkin et al.
-(2001); Gong et al.
-(2024).
-Owing to its computational advantages, circular convolution has also been explored in machine learning for generating long embeddings of high-dimensional data Yu et al.
-(2014) and compressing heavily parameterized layers Cheng et al.
-(2015); Ding et al.
-(2017).
-Remarkably, it achieves these efficiencies without significant performance degradation, which makes it a promising technique for fine-tuning applications.
-Despite its success in small neural networks such as LeNet Cheng et al.
-(2015), circular convolution has not demonstrated lossless performance in modern large foundational models (LFMs) or even in their base architecture, the transformer.
-This limitation may be attributed to the conflict between its high intrinsic bias (i.e., the circulant pattern) and the vast amount of data required for training LFMs.
-Conversely, when fine-tuning LFMs, it is often impractical to collect as much data as needed for training from scratch.
-In such scenarios, the intrinsic bias of circular convolution could potentially serve as a regularization mechanism, thereby benefiting the optimization process of fine-tuning.
-3 Method
+Here, $b$ is the block size and it should be a common divisor (CD) of $d_{1}$ and $d_{2}$.
 In this section, we present C3A (see an overview in Figure 2), a novel PEFT approach based on the circular convolution.
@@
 However, instead of using low-rank decomposition and the matrix multiplication operator, C3A resorts to circular convolution as this additive linear operation.
-Section 3.1 introduces the notations we use.
-Section 3.2 discusses the circular convolution operator, its equivalent circulant matrix, and its calculation in the frequency domain.
-Section 3.3 details an efficient method for backpropagation.
-Section 3.4 describes block-wise convolution for controlling the number of trainable parameters.
-Finally, Section 3.5 analyzes the computational complexity.
 3.1 Notations
@@
 The loss function is represented by $\mathcal{L}$.
-The Fast Fourier Transform and its inverse are denoted by $\mathrm{FFT}$ and $\mathrm{iFFT}$, respectively.
+The Fast Fourier Transform and its inverse are denoted by $\mathrm{FFT}$ and $\mathrm{iFFT}$.
 The Hadamard product is denoted by $\circ$.
@@
 $\mathcal{C}(\Delta\mathbf{w})=\begin{bmatrix}\Delta w_{1}&\Delta w_{2}&\cdots&\Delta w_{d-1}&\Delta w_{d}\\ \Delta w_{d}&\Delta w_{1}&\cdots&\Delta w_{d-2}&\Delta w_{d-1}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ \Delta w_{3}&\Delta w_{4}&\cdots&\Delta w_{1}&\Delta w_{2}\\ \Delta w_{2}&\Delta w_{3}&\cdots&\Delta w_{d}&\Delta w_{1}\\ \end{bmatrix}.$
-Theoretically, the rank of $\mathcal{C}(\Delta\mathbf{w})$ is given by $d-\mathrm{Deg}(\mathrm{gcd}(f(x),x^{d}-1))$ Ingleton (1956), where $\mathrm{Deg}(\cdot)$ denotes the degree of a polynomial, $f(x)$ is the polynomial associated with $\Delta\mathbf{w}$ (i.e., $f(x)=\sum_{i=1}^{d}\Delta w_{i}x^{i-1}$), and $\mathrm{gcd}(\cdot)$ represents the greatest common divisor.
+Theoretically, the rank of $\mathcal{C}(\Delta\mathbf{w})$ is given by $d-\mathrm{Deg}(\mathrm{gcd}(f(x),x^{d}-1))$ (Ingleton, 1956), where $\mathrm{Deg}(\cdot)$ denotes the degree of a polynomial, $f(x)$ is the polynomial associated with $\Delta\mathbf{w}$ (i.e., $f(x)=\sum_{i=1}^{d}\Delta w_{i}x^{i-1}$), and $\mathrm{gcd}(\cdot)$ represents the greatest common divisor.
 Consequently, the theoretical upper bound on the rank of $\mathcal{C}(\Delta\mathbf{w})$ is $d$.
 By learning $\Delta\mathbf{w}$ in the $\mathbb{R}^{n}$ oracle, C3A automatically achieves dynamic rank selection, which is not linearly constrained by the number of learnable parameters, unlike LoRA.
-To achieve high efficiency, we leverage the beautiful circulant structure of $\mathcal{C}(\Delta\mathbf{w})$, which makes it diagonalizable by the Fourier basis ($\mathbf{F}$) .
-In math, it can be described as $\mathcal{C}(\Delta\mathbf{w})=\mathbf{F}\frac{\Lambda}{d}\mathbf{F}^{-1}$ Golub & Van Loan (1996), where $\Lambda$ is its eigenvalues and can be calculated by a Fourier transform of the first row (i.e., $\Lambda=\mathrm{diag}(\mathbf{F}\Delta\mathbf{w})$).
+To achieve high efficiency, enlightened by Ding et al.
+(2017), we leverage the beautiful circulant structure of $\mathcal{C}(\Delta\mathbf{w})$, which makes it diagonalizable by the Fourier basis ($\mathbf{F}$) .
+In math, it can be described as $\mathcal{C}(\Delta\mathbf{w})=\mathbf{F}\frac{\Lambda}{d}\mathbf{F}^{-1}$ (Golub and Van Loan, 1996), where $\Lambda$ is its eigenvalues and can be calculated by a Fourier transform of the first row (i.e., $\Lambda=\mathrm{diag}(\mathbf{F}\Delta\mathbf{w})$).
 Therefore, we can calculate $\Delta\mathbf{w}\star\mathbf{x}$ as
@@
 3.3 Backpropagation
-To implement backpropagation efficiently, we require the analytical derivatives of the loss function $\mathcal{L}$, with respect to $\Delta\mathbf{w}$ and $\mathbf{x}$.
-Utilizing the chain rule, these derivatives can be expressed as:
+To effectuate backpropagation with optimal efficiency, it is imperative to obtain the analytical derivatives of the loss function $\mathcal{L}$ with respect to $\Delta\mathbf{w}$ and $\mathbf{x}$.
+Following the approach outlined in Ding et al.
+(2017), we aim to explain backpropagation using simpler language.
+By applying the chain rule, these derivatives are delineated as follows:
 $\frac{\partial\mathcal{L}}{\partial\mathbf{x}}=\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}},\qquad\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{w}}=\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}}.$ (2)
-Since $\Delta\mathbf{z}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, it follows that $\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})$.
-For $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}$, we note that the circular convolution operation is commutative (i.e., $\mathcal{C}(\Delta\mathbf{w})\mathbf{x}=\mathcal{C}(\mathbf{x})\Delta\mathbf{w}$), thus $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})$.
-Substituting these results into Equation 2, we obtain:
+Given that $\Delta\mathbf{z}=\mathcal{C}(\Delta\mathbf{w})\mathbf{x}$, it logically follows that $\frac{\partial\Delta\mathbf{z}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})$.
+Concerning $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}$, we observe the commutative property of the circular convolution operation (i.e., $\mathcal{C}(\Delta\mathbf{w})\mathbf{x}=\mathcal{C}(\mathbf{x})\Delta\mathbf{w}$), which implies $\frac{\partial\Delta\mathbf{z}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})$.
+Substituting these findings into Equation 2, we derive:
 $\frac{\partial\mathcal{L}}{\partial\mathbf{x}}=\mathcal{C}(\Delta\mathbf{w})\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}},\qquad\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{w}}=\mathcal{C}(\mathbf{x})\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}}.$
@@
 $\frac{\partial\mathcal{L}}{\partial\mathbf{x}}=\Delta\mathbf{w}\star\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}},\qquad\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{w}}=\mathbf{x}\star\frac{\partial\mathcal{L}}{\partial\Delta\mathbf{z}}.$
-By explicitly implementing this derivative calculation using Equation 1, backpropagation can also enjoy the computational efficiency provided by the FFT algorithm.
+By meticulously executing this derivative computation in accordance with Equation 1, backpropagation can harness the computational efficacy facilitated by the FFT algorithm.
 3.4 Block-Circular Convolution
-Despite the elegance and efficiency of the circular convolution operator, it has 2 fundamental limitations due to the requirement that the convolution kernel must be the same size as the activation vector: ❶ It cannot be applied to non-square weight matrices.The number of learnable parameters cannot be altered.
-The first limitation restricts its application in scenarios such as fine-tuning a LLaMA3-8B model, where the weight matrix sizes include $4096\times 1024$.
-The second limitation reduces the flexibility of C3A, making it challenging to handle complex downstream tasks that require a larger number of learnable parameters.
-To address these limitations, we resort to block-circular convolution.
-By splitting the activation vector $\mathbf{x}$ and the post-transformation vector $\mathbf{z}$ into blocks of the same size, we can assign unique convolution kernels to each pair of these blocks.
-To be concrete,
-$\mathbf{x}=\begin{bmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\cdots&\mathbf{x}_{\frac{d_{2}}{b}}\end{bmatrix},\qquad\Delta\mathbf{z}=\begin{bmatrix}\Delta\mathbf{z}_{1}&\Delta\mathbf{z}_{2}&\cdots&\Delta\mathbf{z}_{\frac{d_{1}}{b}}\end{bmatrix},$
+Notwithstanding the elegance and efficiency of the circular convolution operator, it is subject to two fundamental limitations stemming from the constraint that the convolution kernel must match the dimensions of the activation vector: ① It is inapplicable to non-square weight matrices.The count of learnable parameters remains fixed.
+The first restriction hampers its applicability in scenarios such as fine-tuning a LLaMA3-8B model, where the weight matrix dimensions include $4096\times 1024$.
+The second constraint diminishes the adaptability of C3A, presenting challenges in addressing complex downstream tasks that necessitate a greater number of learnable parameters.
+To mitigate these limitations, we employ block-circular convolution (Ding et al., 2017).
+By partitioning the activation vector $\mathbf{x}$ and the post-transformation vector $\Delta\mathbf{z}$ into blocks of identical size, unique convolution kernels can be allocated to each pair of these blocks.
+Specifically,
+$\mathbf{x}$ $=\begin{bmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\cdots&\mathbf{x}_{\frac{d_{2}}{b}}\end{bmatrix}$
+$\Delta\mathbf{z}$ $=\begin{bmatrix}\Delta\mathbf{z}_{1}&\Delta\mathbf{z}_{2}&\cdots&\Delta\mathbf{z}_{\frac{d_{1}}{b}}\end{bmatrix},$
 where $b$ is the block size and $b$ need to be a common divisor of $d_{1}$ and $d_{2}$.
@@
 This calculation can be represented by a block-circular matrix:
-$\Delta\mathbf{z}=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{x},\qquad\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})=\begin{bmatrix}\mathcal{C}(\Delta\mathbf{w}_{11})&\mathcal{C}(\Delta\mathbf{w}_{12})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{1\frac{d_{2}}{b}})\\ \mathcal{C}(\Delta\mathbf{w}_{21})&\mathcal{C}(\Delta\mathbf{w}_{22})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{2\frac{d_{2}}{b}})\\ \cdots&\cdots&\cdots&\cdots\\ \mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}1})&\mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}2})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}\frac{d_{2}}{b}})\end{bmatrix}.$
-In this case, $\Delta\mathbf{w}_{ij}\in\mathbb{R}^{b}$ and we have $\frac{d_{1}d_{2}}{b^{2}}b=\frac{d_{1}d_{2}}{b}$ learnable parameters.
-In fact, $b$ acts as a hyperparameter controlling the number of learnable parameters, just like $r$ for LoRA.
-However, we highlight that while $r$ controls the rank of the delta matrix and learnable parameter number at the same time, $b$ only affects the latter.
-Disentangling these attributes enables higher adaptation flexibility and potential better results.
+$\Delta\mathbf{z}$ $=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{x}$ (3)
+$\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})$ $=\begin{bmatrix}\mathcal{C}(\Delta\mathbf{w}_{11})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{1\frac{d_{2}}{b}})\\ \mathcal{C}(\Delta\mathbf{w}_{21})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{2\frac{d_{2}}{b}})\\ \cdots&\cdots&\cdots\\ \mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}1})&\cdots&\mathcal{C}(\Delta\mathbf{w}_{\frac{d_{1}}{b}\frac{d_{2}}{b}})\end{bmatrix}.$ (4)
+We refer our readers to Algorithm A1 in Appendix A for a Pytorch implementation.
+In this context, $\Delta\mathbf{w}_{ij}\in\mathbb{R}^{b}$, and it follows that $\frac{d_{1}d_{2}}{b^{2}}b=\frac{d_{1}d_{2}}{b}$ represents the number of learnable parameters.
+Notably, the parameter $b$ serves as a hyperparameter modulating the quantity of learnable parameters, analogous to the role of $r$ in LoRA.
+It is imperative to distinguish, however, that whereas $r$ simultaneously governs the rank of the delta matrix and the number of learnable parameters, $b$ exclusively influences the latter.
+This disentanglement of matrix rank and parameter count facilitates greater adaptability and potentially yields superior outcomes.
 3.5 Complexity Analysis
@@
 3.5.1 Time Complexity
-LoRA incorporates low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, which are sequentially multiplied with the activation vector, resulting in a time complexity of $\mathcal{O}(r(d_{1}+d_{2}))$.
-Typically, $r\ll\min(d_{1},d_{2})$.
-Conversely, VeRA, despite its high-rank nature and minimal number of trainable parameters, incurs a prohibitive time complexity of $\mathcal{O}(r_{v}(d_{1}+d_{2}))$, where $r_{v}$ can exceed $\max(d_{1},d_{2})$.
-Therefore, striking a balance between high rank and computational efficiency is never a low-hanging fruit.
-Thanks to the $\mathcal{O}(n\log n)$ complexity of the FFT algorithm used in Equation 1, C3A achieves a time complexity of $\mathcal{O}((d_{1}+d_{2})\log{b}+\frac{d_{1}d_{2}}{b})$.
-The first term is the time complexity for FFT, and the second term is for aggregation.
+LoRA integrates low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, which are successively multiplied with the activation vector, resulting in a computational complexity of $\mathcal{O}(r(d_{1}+d_{2}))$.
+Generally, $r\ll\min(d_{1},d_{2})$.
+In contrast, VeRA, despite its high-rank structure and relatively few trainable parameters, suffers from a prohibitive computational complexity of $\mathcal{O}(r_{v}(d_{1}+d_{2}))$, where $r_{v}$ can exceed $\max(d_{1},d_{2})$.
+Consequently, striking an optimal balance between high rank and computational efficiency remains an elusive task.
+On GPUs, the cuFFT backend automatically parallelizes FFT operations along the axes not being transformed, with the degree of parallelism $p$ determined by the available resources.
+Thanks to the $\mathcal{O}(n\log n)$ complexity of the FFT algorithm used in Equation 1, C3A achieves a time complexity of $\mathcal{O}(\frac{(d_{1}+d_{2})}{p}\log{b}+\frac{d_{1}d_{2}}{b})$.
+The first term is the time complexity for FFT and the second term is for aggregation.
 In practical scenarios, $b$ is chosen as the greatest common divisor of $d_{1}$ and $d_{2}$ to achieve a high compression ratio.
-Given that, C3A is comparable to LoRA in terms of time complexity.
+Given that, C3A is comparable to LoRA in time complexity.
+Method Time Space
+# Param # Other
+LoRA $\mathcal{O}(r(d_{1}+d_{2}))$ $r(d_{1}+d_{2})$ 0
+VeRA $\mathcal{O}(r_{v}(d_{1}+d_{2}))$ $r_{v}+d_{1}$ $r_{v}(d_{1}+d_{2})$
+C3A $\mathcal{O}(\frac{d_{1}+d_{2}}{p}\log{b}+\frac{d_{1}d_{2}}{b})$ $\frac{d_{1}d_{2}}{b}$ $pb$
 Table 1: Time and space complexity comparison of LoRA, VeRA and C3A.
@@
 We highlight that in practice, to achieve similar performance, $\frac{\max(d_{1},d_{2})}{b}\leq r\ll r_{v}$.
-Method Time # Param # Other # Total
-LoRA $\mathcal{O}(r(d_{1}+d_{2}))$ $r(d_{1}+d_{2})$ 0 $r(d_{1}+d_{2})$
-VeRA $\mathcal{O}(r_{v}(d_{1}+d_{2}))$ $r_{v}+d_{1}$ $r_{v}(d_{1}+d_{2})$ $r_{v}(d_{1}+d_{2})+r_{v}+d_{1}$
-C3A $\mathcal{O}((d_{1}+d_{2})\log{b}+\frac{d_{1}d_{2}}{b})$ $\frac{d_{1}d_{2}}{b}$ $b$ $\frac{d_{1}d_{2}}{b}+b$
 3.5.2 Space Complexity
@@
 LoRA does not rely on auxiliary tensors, while VeRA necessitates 2 random projection matrices, with a total size of $r_{v}(d_{1}+d_{2})$.
-Since $r_{v}$ is by no means negligible, the memory consumption of VeRA would be significantly larger than that of LoRA.
-In terms of C3A, the only additional auxiliary tensor would be of size $b\leq\min(d_{1},d_{2})$, which is reserved by the FFT algorithm.
-Given a proper $b$, which is often close to the greatest common divisor of $d_{1}$ and $d_{2}$, the space complexity of C3A excels.
-Additionally, we observe that when applying FFT, input blocks and output blocks are independent within their own group.
-Therefore, parallelization can be employed to enable the flexibility of trading space for time.
+Since $r_{v}$ is by no means negligible, the memory usage of VeRA is significantly larger than that of LoRA.
+In terms of C3A, the only additional auxiliary tensor would be of size $pb\leq\min(d_{1},d_{2})$, which is reserved by the FFT algorithm.
+By selecting an appropriate $b$, which is often close to the greatest common divisor of $d_{1}$ and $d_{2}$, the space complexity of C3A is optimized.
+Furthermore, because $p$ scales with the available resources, the algorithm inherently manages dynamic memory consumption without additional effort.
 4 Experiment
-In this section, we firstly conduct experiments on a synthetic dataset to show the superior expressiveness of C3A over LoRA.
-After that, we comprehensively evaluate C3A in both natural language processing (NLP) and computer vision (CV).
-For NLP, we first demonstrate the effectiveness of C3A using the GLUE benchmark with RoBERTa-Base and RoBERTa-Large models.
-Additionally, we apply C3A to fine-tuning the LLaMA family models, which stands as a representative of prevalent large language models (LLMs).
-For CV, we examine the most representative classification task using Vision Transformers (ViTs) on a wide range of datasets.
-4.1 Synthetic Data
-Settings.
-We distribute 8 points evenly on a 2D plane as cluster centers and randomly sample 30 points from 8 corresponding multivariate Gaussian distributions.
-A 3-layer MLP is then used to classify these point clusters.
-To compare the expressiveness of 2 types of layers, we replace the middle layer with either a low-rank layer or a circulant layer, ensuring that both layers have the same number of trainable parameters for a fair comparison.
-Figure 3: Expressiveness test on synthetic data.
-The left figure shows the synthetic data used for the experiment, while the right figure illustrates the training accuracy curves of a 3-layer MLP, incorporating C3A, LoRA, and standard linear layers, respectively.
-Results.
-The results are presented in Figure 3.
-We observe that LoRAr=1 struggles with this straightforward classification task.
-In contrast, C3Ab=128/2, despite using the same number of parameters, achieves perfect classification, comparable to a standard linear layer.
-This demonstrates the high expressiveness of C3A given the same parameter budget.
-4.2 Natural Language Understanding
-Baselines.
-We compare our C3A with several representative PEFT methods, including BitFit Zaken et al.
-(2021), (IA)3 Liu et al.
-(2022), LoRA Hu et al.
-(2021), VeRA Kopiczko et al.
-(2023), and BOFT Liu et al.
-(2023).
-BitFit selectively fine-tunes existing parameters, specifically the biases.
-(IA)3 is the state-of-the-art method that adds extra adapters.
-LoRA is a widely known PEFT method that employs low-rank decomposition to compress additive delta matrices.
-VeRA is a recent approach that focuses on further reducing trainable parameters of LoRA while preserving a high rank.
-BOFT is another innovative method in PEFT research, compressing multiplicative delta matrices using orthogonal decomposition and butterfly factorization.
+Methods # Params Mem SST-2 MRPC CoLA QNLI RTE STS-B Avg.
+Base Full 124M 17.19G 94.01$\pm$0.39 87.10$\pm$0.79 62.00$\pm$1.16 92.40$\pm$0.28 77.33$\pm$2.68 90.70$\pm$0.14 83.92
+BitFit 0.102M 12.60G 93.30$\pm$0.30 85.80$\pm$0.21 59.21$\pm$1.74 91.96$\pm$0.18 73.07$\pm$1.34 90.18$\pm$0.17 82.25
+(IA)3 0.111M 19.86G 92.98$\pm$0.34 85.86$\pm$0.59 60.49$\pm$1.09 91.56$\pm$0.17 69.10$\pm$1.18 90.06$\pm$0.21 81.67
+LoRAr=8 0.295M 13.75G 94.50$\pm$0.41 85.68$\pm$0.74 60.95$\pm$1.57 92.54$\pm$0.20 76.68$\pm$1.42 89.76$\pm$0.39 83.35
+VeRAr=1024 0.043M 15.51G 93.97$\pm$0.17 86.23$\pm$0.41 62.24$\pm$1.91 91.85$\pm$0.17 75.74$\pm$1.56 90.27$\pm$0.25 83.38
+[…]
+C3Ab=768/1 0.018M 12.83G 93.42$\pm$0.26 86.33$\pm$0.32 61.83$\pm$0.96 91.83$\pm$0.04 76.17$\pm$1.39 90.46$\pm$0.29 83.34
+C3Ab=768/6 0.111M 12.72G 94.20$\pm$0.16 86.67$\pm$0.54 62.48$\pm$1.20 92.32$\pm$0.25 77.18$\pm$1.41 90.16$\pm$0.42 83.84
+Large Full 354M 43.40G 95.75$\pm$0.45 88.35$\pm$0.64 64.87$\pm$1.25 92.40$\pm$0.28 84.48$\pm$1.14 91.65$\pm$0.14 86.25
+BitFit 0.271M 30.65G 95.09$\pm$0.27 88.10$\pm$0.76 65.40$\pm$0.76 94.06$\pm$0.14 82.60$\pm$1.15 91.73$\pm$0.20 86.16
+(IA)3 0.295M 48.81G 95.32$\pm$0.20 87.06$\pm$0.57 66.52$\pm$1.10 94.18$\pm$0.15 84.33$\pm$2.38 91.58$\pm$0.39 86.50
+LoRAr=8 0.786M 34.12G 95.53$\pm$0.35 86.12$\pm$0.86 65.16$\pm$0.76 93.73$\pm$0.30 83.75$\pm$0.51 91.46$\pm$0.21 85.96
+VeRAr=256 0.061M 34.16G 95.83$\pm$0.43 87.72$\pm$0.55 63.66$\pm$1.45 94.11$\pm$0.20 83.03$\pm$1.65 91.12$\pm$0.37 85.91
+[…]
+C3Ab=1024/1 0.049M 31.83G 95.78$\pm$0.05 88.02$\pm$0.62 66.59$\pm$1.20 94.22$\pm$0.25 82.89$\pm$0.67 91.86$\pm$0.14 86.56
+C3Ab=1024/8 0.393M 31.79G 95.78$\pm$0.15 88.09$\pm$0.47 67.18$\pm$1.92 94.26$\pm$0.19 84.62$\pm$1.36 91.81$\pm$0.36 86.96
 Table 2: Performance of different PEFT methods on the GLUE benchmark.
@@
 The best results for each dataset are highlighted in bold.
-# Trainable parameters does not include the classification head since each method uses a head of the same size.
-Memory Cost is measured on fixed length (i.e., $256$) data with a batchsize of $64$.
-Method # Trainable Parameters Memory Cost (GB) SST-2 MRPC CoLA QNLI RTE STS-B Avg.
-Base Full 124M 17.19 94.01$\pm$0.39 87.10$\pm$0.79 62.00$\pm$1.16 92.40$\pm$0.28 77.33$\pm$2.68 90.70$\pm$0.14 83.92
-BitFit 102K 12.60 93.30$\pm$0.30 85.80$\pm$0.21 59.21$\pm$1.74 91.96$\pm$0.18 73.07$\pm$1.34 90.18$\pm$0.17 82.25
-(IA)3  [^1] K 19.86 92.98$\pm$0.34 85.86$\pm$0.59 60.49$\pm$1.09 91.56$\pm$0.17 69.10$\pm$1.18 90.06$\pm$0.21 81.67
-LoRAr=8 295K 13.75 94.50$\pm$0.41 85.68$\pm$0.74 60.95$\pm$1.57 92.54$\pm$0.20 76.68$\pm$1.42 89.76$\pm$0.39 83.35
-VeRAr=1024 43K 15.51 93.97$\pm$0.17 86.23$\pm$0.41 62.24$\pm$1.91 91.85$\pm$0.17 75.74$\pm$1.56 90.27$\pm$0.25 83.38
+# Params does not include the classification head since each method uses a head of the same size.
+Memory cost (Mem) is measured on fixed length (i.e., $256$) data with a batchsize of $64$.
+Baselines.
+We compare our C3A with several representative PEFT methods, including BitFit (Zaken et al., 2021), (IA)3 (Liu et al., 2022), LoRA (Hu et al., 2021), VeRA (Kopiczko et al., 2023), BOFT (Liu et al., 2023) and DoRA (Liu et al., 2024b).
+BitFit fine-tunes biases.
+(IA)3 is a SOTA method adding adapters.
+LoRA uses low-rank decomposition to compress additive delta matrices.
+VeRA reduces LoRA’s trainable parameters while maintaining a high rank.
+BOFT compresses multiplicative delta matrices via orthogonal decomposition and butterfly factorization.
+DoRA separately learns magnitude and direction of delta matrices.
+Methods Params (%) Mem BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Avg.
+ChatGPT - - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
+LLaMA2-7B LoRAr=32 0.83 41.71G 71.0 81.4 79.6 87.4 83.2 82.6 67.5 81.5 79.3
+VeRAr=16384 0.05 57.88G 68.9 $\downarrow 2.1$ 81.0 $\downarrow 0.4$ 77.4 $\downarrow 2.2$ 87.5 $\uparrow 0.1$ 81.3 $\downarrow 1.9$ 81.1 $\downarrow 1.5$ 62.8 $\downarrow 4.7$ 79.9 $\downarrow 1.6$ 77.5 $\downarrow 1.8$
 […]
-C3Ab=768/1 18K 12.83 93.42$\pm$0.26 86.33$\pm$0.32 61.83$\pm$0.96 91.83$\pm$0.04 76.17$\pm$1.39 90.46$\pm$0.29 83.34
-C3Ab=768/6  [^1] K 12.72 94.20$\pm$0.16 86.67$\pm$0.54 62.48$\pm$1.20 92.32$\pm$0.25 77.18$\pm$1.41 90.16$\pm$0.42 83.84
-Large BitFit 271K 30.65 95.09$\pm$0.27 88.10$\pm$0.76 65.40$\pm$0.76 94.06$\pm$0.14 82.60$\pm$1.15 91.73$\pm$0.20 86.16
-(IA)3 295K 48.81 95.32$\pm$0.20 87.06$\pm$0.57 66.52$\pm$1.10 94.18$\pm$0.15 84.33$\pm$2.38 91.58$\pm$0.39 86.50
-LoRAr=8 786K 34.12 95.53$\pm$0.35 86.12$\pm$3.22 65.16$\pm$0.76 93.73$\pm$0.30 83.75$\pm$0.51 91.46$\pm$0.21 85.96
-VeRAr=256 61K 34.16 95.83$\pm$0.43 87.72$\pm$0.55 63.66$\pm$1.45 94.11$\pm$0.20 83.03$\pm$1.65 91.12$\pm$0.37 85.91
+DoRAr=32 0.84 55.24G 72.3 $\uparrow 1.3$ 84.2 $\uparrow 2.8$ 78.7 $\downarrow 0.9$ 88.5 $\uparrow 1.1$ 84.0 $\uparrow 0.8$ 80.7 $\downarrow 1.9$ 69.1 $\uparrow 1.6$ 83.7 $\uparrow 2.2$ 80.2 $\uparrow 0.9$
+C3Ab=4096/32 0.35 44.89G 73.4 $\uparrow 2.4$ 83.4 $\uparrow 2.0$ 82.1 $\uparrow 2.5$ 88.3 $\uparrow 0.9$ 84.9 $\uparrow 1.7$ 86.6 $\uparrow 4.0$ 66.7 $\downarrow 0.8$ 82.5 $\uparrow 1.0$ 81.0 $\uparrow 1.7$
+LLaMA3-8B LoRAr=32 0.70 51.18G 73.8 88.2 80.4 94.0 85.5 87.5 78.1 84.0 83.9
+VeRAr=16384 0.04 66.03G 72.2 $\downarrow 1.6$ 87.5 $\downarrow 0.7$ 80.0 $\downarrow 0.4$ 94.3 $\uparrow 0.3$ 83.9 $\downarrow 1.6$ 84.5 $\downarrow 3.0$ 75.6 $\downarrow 2.5$ 82.4 $\downarrow 1.6$ 82.5 $\downarrow 1.4$
 […]
-C3Ab=1024/1 49K 31.83 95.78$\pm$0.05 88.02$\pm$0.62 66.59$\pm$1.20 94.22$\pm$0.25 82.89$\pm$0.67 91.86$\pm$0.14 86.56
-C3Ab=1024/8 393K 31.79 95.78$\pm$0.15 88.09$\pm$0.47 67.18$\pm$1.92 94.26$\pm$0.19 84.62$\pm$1.36 91.81$\pm$0.36 86.96
+DoRAr=32 0.71 63.37G 75.3 $\uparrow 1.5$ 88.9 $\uparrow 0.7$ 82.2 $\uparrow 1.8$ 96.7 $\uparrow 2.7$ 86.0 $\uparrow 0.5$ 89.0 $\uparrow 1.5$ 79.8 $\uparrow 1.7$ 84.9 $\uparrow 0.9$ 85.3 $\uparrow 1.4$
+C3Ab=4096/32 0.26 56.08G 76.9 $\uparrow 3.1$ 91.4 $\uparrow 3.2$ 82.1 $\uparrow 1.7$ 94.9 $\uparrow 0.9$ 86.9 $\uparrow 1.4$ 89.6 $\uparrow 2.1$ 79.4 $\uparrow 1.3$ 86.1 $\uparrow 2.1$ 85.9 $\uparrow 2.0$
+Table 3: Comparison of various methods on the LLaMA2-7B and LLaMA3-8B models across eight commonsense reasoning datasets.
+For each language model, the best result on each dataset is shown in bold.
+The symbols $\uparrow$ and $\downarrow$ indicate relative improvement and decrease, respectively, compared to the LoRA method.
+Experiments are conducted on a single H800 GPU with 80GB HBM.
+Methods Params (%) Math Avg. Code Avg.
+GSM8K MATH HumanEval HumanEval+ MBPP MBPP+
+LLaMA2-7B Zero-Shot - 7.3 1.1 4.2 11.0 9.8 30.2 24.1 18.8
+LoRAr=32 0.83 60.5 11.7 36.1 31.7 28.0 35.4 30.4 31.4
+VeRAr=16384 0.05 58.5 $\downarrow 2.0$ 9.4 $\downarrow 2.3$ 34.0 $\downarrow 2.1$ 28.9 $\downarrow 2.8$ 24.6 $\downarrow 3.4$ 35.1 $\downarrow 0.3$ 29.6 $\downarrow 0.8$ 29.5 $\downarrow 1.9$
+DoRAr=32 0.84 62.3 $\uparrow 1.8$ 12.4 $\uparrow 0.7$ 37.3 $\uparrow 1.2$ 33.2 $\uparrow 1.5$ 28.8 $\uparrow 0.8$ 36.8 $\uparrow 1.4$ 31.8 $\uparrow 1.4$ 32.7 $\uparrow 1.3$
+C3Ab=4096/32 0.35 61.5 $\uparrow 1.0$ 13.0 $\uparrow 1.3$ 37.2 $\uparrow 1.1$ 33.8 $\uparrow 2.1$ 29.0 $\uparrow 1.0$ 36.5 $\uparrow 1.1$ 31.9 $\uparrow 1.5$ 32.8 $\uparrow 1.4$
+LLaMA3-8B Zero-Shot - 33.1 5.3 19.2 33.5 29.3 61.4 51.6 44.0
+LoRAr=32 0.70 77.2 28.2 52.7 57.9 52.4 64.8 55.3 57.6
+VeRAr=16384 0.04 76.0 $\downarrow 1.2$ 26.7 $\downarrow 1.5$ 51.4 $\downarrow 1.3$ 56.3 $\downarrow 1.6$ 50.8 $\downarrow 1.6$ 64.0 $\downarrow 0.8$ 53.6 $\downarrow 1.7$ 56.2 $\downarrow 1.4$
+DoRAr=32 0.71 77.2 $\uparrow 0.0$ 30.0 $\uparrow 1.8$ 53.6 $\uparrow 0.9$ 59.1 $\uparrow 1.2$ 52.9 $\uparrow 0.5$ 65.7 $\uparrow 0.9$ 55.8 $\uparrow 0.5$ 58.4 $\uparrow 0.8$
+C3Ab=4096/32 0.26 78.4 $\uparrow 1.2$ 29.9 $\uparrow 1.7$ 54.1 $\uparrow 1.4$ 59.5 $\uparrow 1.6$ 53.8 $\uparrow 1.4$ 65.4 $\uparrow 0.6$ 56.1 $\uparrow 0.8$ 58.7 $\uparrow 1.1$
+Table 4: Comparison of various methods on the LLaMA2-7B and LLaMA3-8B models across math reasoning and code generation datasets.
+For each language model, the best result on each dataset is shown in bold.
+The symbols $\uparrow$ and $\downarrow$ indicate relative improvement and decrease, respectively, compared to the LoRA method.
+Experiments are conducted on a single H800 GPU with 80GB HBM.
+4.1 GLUE Benchmark
 Settings.
-We evaluate our proposed C3A on the General Language Understanding Evaluation (GLUE) benchmark Wang et al.
-(2018), which encompasses a wide range of natural language understanding (NLU) tasks, including single-sentence classification, similarity and paraphrase, and natural language inference.
-To enhance practicality, we split these datasets following the train-validation-test approach.
-The best-performing model is selected based on validation set performance across the fine-tuning epochs, and the reported performance corresponds to its performance on the test set.
-For this evaluation, we fine-tune the pre-trained RoBERTa-Base and RoBERTa-Large models Liu et al.
+We assess our C3A on the GLUE benchmark (Wang et al., 2018) covering various tasks like single-sentence classification, similarity, paraphrase, and inference.
+See Table A3 in Appendix C for more details.
+Datasets are split into train, validation, and test sets.
+Models are chosen based on validation performance and evaluated on the test set.
+We fine-tune RoBERTa-Base and RoBERTa-Large models (Liu et al., 2019).
+Unique hyperparameters are from original papers (e.g., for VeRA’s $r$ and BOFT’s $b$ and $m$).
+Count of trainable parameters (# Params) exclude the classification head, which is uniform across methods.
+Shared hyperparameters (i.e., learning rates for the classification head and other parameters) are determined by hyperparameter search.
+Regarding the memory test, input data is fixed to 256 tokens with a batch size of 64 for consistency.
+Results.
+Results are presented in Table 2.
+Overall, C3Ab=768/1 and C3Ab=1024/1 achieve superior or comparable performance to baseline methods, despite using an exceptionally small number of trainable parameters.
+As the number of trainable parameters increases, models like C3Ab=768/6 and C3Ab=1024/8 significantly outperform the baselines.
+Moreover, compared to (IA)3, LoRA, VeRA, and BOFT, C3A distinguishes itself with remarkable memory efficiency.
+The only method with better memory efficiency is BitFit, which serves as an upper bound since it does not introduce new parameters.
+Furthermore, most of the delta matrices identified by C3A are of full rank, indicating maximum capacity (Zeng and Lee, 2023) and providing a theoretical basis for outstanding performance.
+4.2 Instruction Fine-tuning
+Commonsense Reasoning.
+Our training utilizes Commonsense170K (Hu et al., 2023), aggregating multiple-choice questions from BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018).
+Evaluation follows Hu et al.
+(2023) and Liu et al.
+(2024b), using greedy search and first-keyword appearance for answer determination.
+Mathematical Reasoning.
+We use MetaMathQA (Yu et al., 2023), containing 395K QA pairs from GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2020).
+Following Yu et al.
+(2023), we evaluate using greedy search on test sets, requiring chain-of-thought reasoning (Wei et al., 2022).
+Code Generation.
+Training uses Magicoder-Evol-Instruct-110k (Wei et al., 2024), a decontaminated subset of WizardCoder (Luo et al., 2024).
+Evaluation on HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and their Plus variants follows EvalPlus (Liu et al., 2024a), reporting Pass@1 metrics.
+Results.
+In Table 3 and Table 4, our principal experimental observations are summarized.
+The C3A framework consistently surpasses LoRA and other baselines within the LLaMA series, with particular efficacy demonstrated in the most recent model, LLaMA3-8B.
+Noteworthy is the significant enhancement in the efficacy of LLaMA3-8B as a foundational model following the implementation of more sophisticated post-training techniques.
+This underscores the criticality of optimizing the fine-tuning protocols for this advanced model.
+It is also remarkable that C3A achieves such results while employing less than half the parameter count of LoRA.
+Furthermore, C3A achieves high rank adaptation and delivers optimal performance while maintaining remarkably low memory consumption.
+In contrast, VeRA and DoRA require substantially greater memory resources, and BoFT results in an OOM condition for a single H800, thereby underscoring the practicality and efficiency of C3A.
+Taken together, the findings robustly underscore the superior efficacy of the C3A methodology.
+We refer readers to Appendix D for examples of model answers after different tuning methods.
+4.3 Ablation Study
+Initialization.
+We investigated initialization effects in C3A compared to LoRA, which is known for initialization sensitivity due to its $\mathbf{A}$ and $\mathbf{B}$ matrices (Hayou et al., 2024a).
+We evaluated multiple initialization strategies across five distinct tasks.
+Each initialization method was tested with 5 independent runs per task, producing a total of 20 runs per task.
+As shown in Figure 3, testing zero, Gaussian, Kaiming uniform, and Xavier uniform initializations for C3A’s convolution kernels revealed performance variations mostly within standard deviations, demonstrating C3A’s robustness to initialization choices.
+Figure 3: Violin plots for runs on different tasks.
+Within each violin plot, a white bar indicates the range of average performance for each initialization strategy.
+It is clear that the choice of initialization does not affect the final result more than the intrinsic stochasticity.
+Expressiveness.
+We demonstrate C3A’s expressiveness on a synthetic dataset by placing 8 cluster centers on a 2D plane and sampling 30 points from their Gaussian distributions.
+A 3-layer MLP is used to classify these clusters.
+To compare the expressiveness of LoRA and C3A, we replace the middle layer with either a low-rank layer or a circulant layer, ensuring that both layers have the same number of trainable parameters for a fair comparison.
+The results are presented in Figure 4.
+We observe that LoRAr=1 struggles with this simple classification task.
+In contrast, C3Ab=128/2, despite using the same number of parameters, achieves a perfect classification, comparable to a standard linear layer.
+This demonstrates the high expressiveness of C3A given the same parameter budget.
+Figure 4: Training curves on the synthetic dataset. We refer our readers to Figure A1 in Appendix E for a visualization of the synthetic dataset.
+Scaling.
+To investigate the scaling effect of C3A, we examine it from both data and model perspectives, following the setup in He et al.
 (2025).
-For the unique hyperparameters of different baselines, we adopt the values suggested in the original papers (e.g., VeRA’s $r$ and BOFT’s $b$ and $m$).
-The number of trainable parameters excludes the classification head, as each method uses one of the same size.
-The shared hyperparameters (i.e., the learning rate for classification head and for other trainable parameters, separately) are found by hyperparameter search.
-For the memory cost, to ensure fairness and consistency, we fix the length of input data to 256 tokens and use a batchsize of 64.
-Results.
-Results are summarized in Table 2.
-In general, C3Ab=768/1 and C3Ab=1024/1 achieves better or on-par performance compared to baseline methods with extremely small trainable parameters.
-When a larger number of trainable parameters is allowed, C3Ab=768/6 and C3Ab=1024/8 significantly outperform baselines.
-In addition, compared to (IA)3, LoRA, VeRA, and BOFT, despite training a similar number of or more parameters, C3A stands out for its excellent memory efficiency.
-The only method with better memory efficiency is BitFit, which can be seen as an upper bound since it does not introduce any new parameters.
-Furthermore, nearly all circulant delta matrices discovered by C3A are full rank, indicating maximal capacity Zeng & Lee (2023) and providing theoretical support for the impressive results.
-4.3 Instruction Tuning
+From a data perspective, we focus on the MATH dataset and vary the training data used.
+As shown in Figure 5, we observe that C3A’s performance improves faster than LoRA as more data is added, suggesting it is more effective at leveraging additional data.
+From the model perspective, we evaluate C3A in small models (LLaMA3-8B, Mistral-7B) and large models (LLaMA3-70B, Mistral-8x7B).
+Results show that C3A outperforms the LoRA baseline in both settings, indicating that its benefits generalize to different model scales.
+Experiments on various data quantities and model sizes showcase C3A’s scalability and robustness, making it a promising approach for adapting language models.
+Figure 5: Data and model scaling of C3A compared to LoRA. LM specifies LLaMA3 and MT specifies Mistral.
+5 Conclusion
+This manuscript introduces C3A, a novel PEFT method.
+Unlike LoRA’s low-rank decomposition, C3A utilizes circular convolution and circulant matrices to construct the delta weight matrix.
+This approach allows independent control over the delta weight matrix’s rank and the number of trainable parameters, enabling high-rank adaptation with limited parameter size.
+By employing FFT in forward and backward propagation, C3A achieves significant computational and memory efficiency, presenting a compelling alternative to LoRA for model fine-tuning.
+6 Limitations
+Despite C3A’s efficiency and effectiveness within limited parameter budgets, it has some intrinsic drawbacks.
+First, due to the requirement of $b\in\mathrm{gcd}(d_{1},d_{2})$, C3A may not be flexible enough to tackle all transformer architectures.
+Besides, the integration of FFT in C3A necessitates specialized kernel support for optimal performance.
+While this support is well-developed and readily available for NVIDIA GPUs, it may not be as mature or optimized for edge devices with different hardware architectures.
+This can potentially hinder the deployment and efficiency of C3A on resource-constrained edge platforms.
+Appendix
+Appendix A Implementations
+Algorithm A1 Block-Circular Convolution PyTorch Implementation
+import torch
+from torch.autograd import Function
+from torch.fft import fft, ifft
+class BlockCircularConvolution(Function):
+@staticmethod
+def forward(ctx, x, w):
+m, n, b = w.shape
+x = x.reshape(*x.shape[:-1], n, b)
+ctx.save_for_backward(x, w)
+x = torch.einsum(
+"...nb,mnb->...mb", ifft(x), fft(w)
+x = fft(x).real
+x = x.reshape(*x.shape[:-2], -1)
+return x
+@staticmethod
+def backward(ctx, grad_output):
+x, w = ctx.saved_tensors
+m, n, b = w.shape
+grad_output = grad_output.reshape(
+*grad_output.shape[:-1], m, b
+grad_output_fft = fft(grad_output)
+x_grad = fft(torch.einsum(
+"...mb,mnb->...nb",
+grad_output_fft, ifft(w)
+)).real
+x_grad = x_grad.reshape(
+*x_grad.shape[:-2], -1
+w_grad = fft(torch.einsum(
+"...mb,...nb->mnb",
+grad_output_fft, ifft(x)
+)).real
+return x_grad, w_grad
+We present the PyTorch implementation of Block-Circular Convolution in Algorithm A1.
+Furthermore, due to the inefficiency of directly assigning entries (as shown in Equation 4), we derive an alternative algorithm to compute the $\Delta\mathbf{W}$ more efficiently.
+Rather than direct assignment, we employ a forward process on the Identity matrix.
+Mathematically, this can be expressed as
+$\Delta\mathbf{W}$ $=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})$
+$=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\cdot\mathbf{I}_{d_{2}}$
+$=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\cdot\left[\mathbf{e}_{1},\mathbf{e}_{2},\cdots,\mathbf{e}_{d_{2}}\right]$
+$=\left[\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{e}_{1},\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{e}_{2},\cdots,\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{e}_{d_{2}}\right]$
+$=\left[\Delta\mathbf{w}\star\mathbf{e}_{1},\Delta\mathbf{w}\star\mathbf{e}_{2},\cdots,\Delta\mathbf{w}\star\mathbf{e}_{d_{2}}\right].$
+Where $\mathbf{I}_{d_{2}}\in\mathbb{R}^{d_{2}\times d_{2}}$ represents an Identity matrix and $\mathbf{e}_{i}$ is the $i$th column of it.
+In pytorch, we can efficiently compute the iFFT of $\{\mathbf{e}_{i}\}_{i=1,2,\cdots,d_{2}}$ by a column-wise iFFT of $\mathbf{I}_{d_{2}}$.
+We present the Pytorch implementation in Algorithm A2 as well.
+Algorithm A2 Fast Algorithm of Getting $\Delta\mathbf{W}$
+import torch
+from torch.fft import fft, ifft
+def get_circulant_fast(w):
+m, n, b = w.shape
+x = torch.eye(n*b)
+x = x.reshape(*x.shape[:-1], n, b)
+x = torch.einsum(
+"...nb,mnb->...mb", ifft(x), fft(w)
+x = fft(x).real.flatten(start_dim=1).T
+return x
+Appendix B Image Classification
 Settings.
-We evaluate C3A against LoRA by fine-tuning LLaMA2-7B/13B and LLaMA3-8B on 4 standard instruction tuning datasets: GSM8k Cobbe et al.
-(2021), MATH Yu et al.
-(2023), ViGGO Juraska et al.
-(2019), and SQL Zhong et al.
-(2017), following the setting outlined in Nikdan et al.
-(2024).
-For the SQL dataset, we preprocess by selecting 12.5% of the dataset, followed by an 4:1 train-test split.
-To ensure a fair comparison, we maintain LoRA parameters with $r=32$ and a dropout rate of 0.05, while exploring various $\alpha$ values ranging from 32 to 128.
-All experiments are conducted for a standard single pass (epoch) over the dataset.
-Results.
-In Table 3, we summarize our main experiments.
-It can be observed that C3A generally surpasses LoRA across both LLaMA2-7B/13B and LLaMA3-8B.
-Notably, C3A only utilizes less than half of the parameter count compared to LoRA.
-Overall, these results clearly highlight the effectiveness of C3A.
-Table 3: Comparison of C3A and LoRA on fine-tuning LLaMA2 and LLaMA3 models in terms of accuracy and trainable parameters.
-The best results for each dataset are highlighted in bold.
-“Avg.” denotes the average accuracy of each method across all datasets.
-Model Method # Trainable Parameters GSM8k MATH ViGGO SQL Avg.
-LLaMA2-7B LoRAr=32 16.8M 27.22 4.10 92.24 65.39 47.24
-C3Ab=4096/32 8.4M 29.34 4.25 94.03 66.21 48.46
-LLaMA2-13B LoRAr=32 26.2M 42.68 6.70 93.19 67.53 52.53
-C3Ab=5120/32 13.1M 42.00 6.65 95.57 67.28 52.88
-LLaMA3-8B LoRAr=32 13.6M 56.57 18.95 91.78 68.19 58.87
-C3Ab=4096/32 5.2M 57.54 19.75 93.07 69.47 59.96
-4.4 Image Classification
-Settings.
-We focus on the task of image classification utilizing the Vision Transformer (ViT) models.
-Specifically, we employ both the Base and Large variants of this prevalent computer vision foundation model, as introduced by Dosovitskiy et al.
-(2020).
-These ViT models are pre-trained on the extensive ImageNet-21K dataset Ridnik et al.
-(2021).
-For the fine-tuning phase, we utilize a diverse set of datasets including Pets, Cars, DTD, EuroSAT, FGVC and RESISC.
-Table 4: Fine-tuning results with ViT-Base and ViT-Large models on various image classification datasets.
+In this study, we concentrate on the task of image classification leveraging Vision Transformer (ViT) models.
+Specifically, we employ both the Base and Large variants of this prominent foundational computer vision model, as delineated by (Dosovitskiy et al., 2020).
+These ViT models undergo pre-training on the expansive ImageNet-21K dataset (Ridnik et al., 2021).
+During the fine-tuning phase, we use an eclectic array of datasets encompassing Pets (Parkhi et al., 2012), Cars (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), FGVC (Maji et al., 2013), and RESISC (Cheng et al., 2017).
+Statistics for these datasets are provided in Table A1.
+Dataset #Train #Validation #Test #Class
+Pets (Parkhi et al., 2012) 3,312 368 3,669 37
+Cars (Krause et al., 2013) 7,329 815 8,041 196
+DTD (Cimpoi et al., 2014) 4,060 452 1,128 47
+EuroSAT (Helber et al., 2019) 16,200 5,400 5,400 10
+FGVC (Maji et al., 2013) 3,000 334 3,333 100
+RESISC (Cheng et al., 2017) 18,900 6,300 6,300 45
+Table A1: Details about the vision datasets.
+Method # Params Pets Cars DTD EuroSAT FGVC RESISC Avg.
+Base Head - 90.28$\pm$0.43 25.76$\pm$0.28 69.77$\pm$0.67 88.72$\pm$0.13 17.44$\pm$0.43 74.22$\pm$0.10 61.03
+Full 85.8M 92.82$\pm$0.54 85.10$\pm$0.21 80.11$\pm$0.56 99.11$\pm$0.07 61.60$\pm$1.00 96.00$\pm$0.23 85.79
+LoRAr=16 0.59M 93.76$\pm$0.44 78.04$\pm$0.33 78.56$\pm$0.62 98.84$\pm$0.08 56.64$\pm$0.55 94.66$\pm$0.17 83.42
+C3Ab=768/12 0.22M 93.88$\pm$0.22 79.05$\pm$0.35 80.57$\pm$0.53 98.88$\pm$0.07 54.31$\pm$0.79 94.54$\pm$0.23 83.54
+Large Head - 91.11$\pm$0.30 37.91$\pm$0.27 73.33$\pm$0.26 92.64$\pm$0.08 24.62$\pm$0.24 82.02$\pm$0.11 66.94
+Full 303M 94.30$\pm$0.31 88.15$\pm$0.50 80.18$\pm$0.66 99.06$\pm$0.10 67.38$\pm$1.06 96.08$\pm$0.20 87.53
+LoRAr=16 1.57M 94.62$\pm$0.47 86.11$\pm$0.42 80.09$\pm$0.42 98.99$\pm$0.03 63.64$\pm$0.83 95.52$\pm$0.21 86.56
+C3Ab=1024/16 0.79M 94.48$\pm$0.30 84.94$\pm$0.39 82.62$\pm$0.52 98.75$\pm$0.19 63.80$\pm$0.37 95.94$\pm$0.16 86.69
+Table A2: Fine-tuning results with ViT-Base and ViT-Large models on various image classification datasets.
 The models are fine-tuned for 10 epochs, and the best-performing model, based on validation set accuracy, is selected.
 The reported accuracy corresponds to the performance on the test set.
-The best results for each dataset are highlighted in bold.
+The best results between LoRA and C3A for each dataset are highlighted in bold.
 “Avg.” denotes the average accuracy of each method across all datasets.
-Method # Trainable Parameters Pets Cars DTD EuroSAT FGVC RESISC Avg.
-Base Head - 90.28$\pm$0.43 25.76$\pm$0.28 69.77$\pm$0.67 88.72$\pm$0.13 17.44$\pm$0.43 74.22$\pm$0.10 61.03
-Full 85.8M 92.82$\pm$0.54 85.10$\pm$0.21 80.11$\pm$0.56 99.11$\pm$0.07 61.60$\pm$1.00 96.00$\pm$0.23 85.79
-LoRAr=16 590K 93.76$\pm$0.44 78.04$\pm$0.33 78.56$\pm$0.62 98.84$\pm$0.08 56.64$\pm$0.55 94.66$\pm$0.17 83.42
-C3Ab=768/12 221K 93.88$\pm$0.22 79.05$\pm$0.35 80.57$\pm$0.53 98.88$\pm$0.07 54.31$\pm$0.79 94.54$\pm$0.23 83.54
-Large Head - 91.11$\pm$0.30 37.91$\pm$0.27 73.33$\pm$0.26 92.64$\pm$0.08 24.62$\pm$0.24 82.02$\pm$0.11 66.94
-Full 303M 94.30$\pm$0.31 88.15$\pm$0.50 80.18$\pm$0.66 99.06$\pm$0.10 67.38$\pm$1.06 96.08$\pm$0.20 87.53
-LoRAr=16 1.57M 94.62$\pm$0.47 86.11$\pm$0.42 80.09$\pm$0.42 98.99$\pm$0.03 63.64$\pm$0.83 95.94$\pm$0.21 86.56
-C3Ab=1024/16 786K 94.48$\pm$0.30 84.94$\pm$0.39 82.62$\pm$0.52 98.75$\pm$0.19 63.80$\pm$0.37 95.52$\pm$0.16 86.69
 Results.
-Table 4 presents a summary of the results obtained from 6 image classification datasets using the ViT Base and Large models.
-The LoRA and C3A methods both show substantial performance improvements over Head Tuning, highlighting their effectiveness in the image classification domain.
-Notably, our method achieves comparable performance to LoRA while utilizing only half of the parameter count.
-5 Conclusion
-In this paper, we introduce C3A, an innovative approach to Parameter-Efficient Fine-Tuning (PEFT).
-Unlike LoRA, which utilizes low-rank decomposition, C3A employs circular convolution and its associated circulant matrix to model the delta weight matrix.
-This approach aims to decouple the rank of the delta weight matrix from the number of learnable parameters, allowing for high-rank adaptation while maintaining a limited parameter size.
-By leveraging the Fast Fourier Transform (FFT) in both the forward and backward passes, C3A achieves significant computational and memory efficiency.
-Consequently, C3A presents itself as a compelling alternative to LoRA for model fine-tuning.
+Table A2 delineates a comprehensive summary of the outcomes derived from six distinct image classification datasets employing the ViT Base and Large models.
+The LoRA and C3A techniques exhibit significant enhancements in performance relative to Head Tuning, thereby underscoring their efficacy within the realm of image classification.
+Remarkably, our methodology demonstrates a performance on par with LoRA while necessitating only half of the parameter count.
+Appendix C GLUE Benchmark Details
+Corpus Task # Train # Val # Test # Labels Metrics Domain
+Single-Sentence Tasks
+CoLA Acceptability 8.55k 1.04k 1.06k 2 Matthews Corr. misc.
+SST-2 Sentiment 67.3k 872 1.82k 2 Accuracy Movie reviews
+Similarity and Paraphrase Tasks
+MRPC Paraphrase 3.67 408 1.73k 2 Accuracy/F1 News
+STS-B Sentence similarity 5.75k 1.5k 1.38k 1 Pearson/Spearman Corr. misc.
+QQP Paraphrase 364k 40.4k 391k 2 Accuracy/F1 Social QA
+Inference Tasks
+MNLI NLI 393k 19.65k 19.65k 3 Accuracy misc.
+QNLI QA/NLI 105k 5.46k 5.46k 2 Accuracy Wikipedia
+RTE NLI 2.49k 277 3k 2 Accuracy News & Wikipedia
+Table A3: Task descriptions and dataset statistics of the GLUE benchmark (Wang et al., 2018).
+The General Language Understanding Evaluation (GLUE) benchmark is a collection of nine natural language understanding tasks designed to test a model’s performance on a diverse set of natural language understanding challenges.
+It was introduced by researchers from New York University, the University of Washington, and DeepMind in a 2018 paper titled “GLUE: A Multi-Task Benchmark and Analysis Platform for Natural Language Understanding” (Wang et al., 2018).
+The nine tasks included in the benchmark are:
+  1. 1.
+    CoLA (Corpus of Linguistic Acceptability): A binary classification task that tests a model’s ability to determine whether a given sentence is grammatically acceptable.
+  2. 2.
+    SST-2 (Stanford Sentiment Treebank): A binary sentiment classification task using movie review excerpts.
+  3. 3.
+    MRPC (Microsoft Research Paraphrase Corpus): A binary classification task to determine whether two sentences are semantically equivalent.
+  4. 4.
+    STS-B (Semantic Textual Similarity Benchmark): A regression task that scores the semantic similarity of sentence pairs on a scale from 1 to 5.
+  5. 5.
+    QQP (Quora Question Pairs): A binary classification task to determine whether two questions are semantically equivalent.
+  6. 6.
+    MNLI (Multi-Genre Natural Language Inference): A ternary classification task that tests a model’s ability to perform textual entailment across multiple genres.
+  7. 7.
+    QNLI (Question Natural Language Inference): A binary classification task that tests a model’s ability to determine whether a sentence contains the answer to a given question.
+  8. 8.
+    RTE (Recognizing Textual Entailment): A binary classification task that tests a model’s ability to determine whether a premise entails a hypothesis.
+  9. 9.
+    WNLI (Winograd Natural Language Inference): A binary classification task that tests a model’s ability to resolve ambiguous pronouns in a sentence.
+The GLUE benchmark has become a widely-used standard for evaluating the performance of pre-trained language models and fine-tuned models on a variety of natural language understanding tasks.
+The benchmark provides a single-number metric, the GLUE score, which is an average of the scores on each individual task.
+This allows for easy comparison of different models and architectures.
+We refer readers to Table A3 for more details.
+Appendix D Instruction Fine-tuning Examples
+Example D.1: MATH Below is an instruction that describes a task.
+Write a response that appropriately completes the request.
+Instruction: Find the number of square units bounded by the $x$-axis, the $y$-axis and the line $x+y=6$.
+Answer the above question.
+First think step by step and then answer the final number.
+LoRA Answer: The line $x+y=6$ intersects the $x$-axis at $x=6$ and the $y$-axis at $y=6$.
+So, the line divides the square with side length 6 into two triangles.
+The area of each triangle is $\frac{1}{2}\times 6\times 6=18$.
+So, the area of the square is $2\times 18=\boxed{36}$.
+The final answer is: 36.
+C3A Answer: The line $x+y=6$ intersects the $x$-axis at $(6,0)$ and the $y$-axis at $(0,6)$.
+Thus, the line divides the first quadrant into two triangles, one with base $6$ and height $6$ and the other with base $6$ and height 00.
+The area of the first triangle is $(6\times 6)/2=18$ square units, and the area of the second triangle is $(6\times 0)/2=0$ square units.
+Thus, the total area is $18+0=\boxed{18}$ square units.
+The final answer is: 18.
+Example D.2: GSM8K Below is an instruction that describes a task.
+Write a response that appropriately completes the request.
+Instruction: Tim makes a special honey and jam mix.
+To make 1 jar of the mixture it takes 2 pounds of passion fruit and 1.5 pounds of jackfruit.
+It also takes 2 pounds of special honey.
+The passion fruit costs \$6 per pound, the jackfruit is \$8 per pound, and the honey is \$10 per pound.
+He sells each jar for \$50.
+How much profit does he make selling 10 jars?
+Answer the above question.
+First think step by step and then answer the final number.
+LoRA Answer: The passion fruit costs $6\times 2=12$.
+The jackfruit costs $8\times 1.5=12$.
+The honey costs $10\times 2=20$.
+So the total cost is $12+12+20=44$.
+So he makes a profit of $50\times 10-44=\boxed{460}$.
+The final answer is: 460.
+C3A Answer: The passion fruit costs $6\times 2=12$.
+The jackfruit costs $8\times 1.5=12$.
+So the fruit costs $12+12=24$.
+The honey costs $10\times 2=20$.
+So the total cost per jar is $24+20=44$.
+So he makes a profit of $50-44=6$ per jar.
+So he makes a profit of $6\times 10=\boxed{60}$.
+The final answer is: 60.
+Appendix E Synthetic Dataset
+We visualize the synthetic dataset we use for the expressiveness ablation study in Figure A1.
+Figure A1: Synthetic dataset used for the expresiveness ablation study.
+Appendix F Hyperparameters
+Model Hyperparameter SST-2 MRPC CoLA QNLI RTE STS-B
+Both Optimizer AdamW
+LR Schedule Linear
+Warmup Ratio 0.06
+C3A Initialization Xavier Uniform
+Max Seq. Len 512
+Base Epochs 40 80 80 40 80 80
+Batch Size 128 128 128 64 64 128
+Learning Rate (C3Ab=768/6) 2E-1 3E-1 2E-1 7E-2 3E-1 2E-1
+Learning Rate (Head) 2E-4 4E-6 3E-2 8E-6 6E-3 4E-2
+Large Epochs 10 80 70 30 60 40
+Batch Size 128 128 128 32 64 128
+Learning Rate (C3Ab=1024/8) 9E-2 3E-1 2E-1 7E-2 5E-2 2E-1
+Learning Rate (Head) 2E-4 5E-6 3E-3 8E-6 3E-3 5E-4
+Table A4: Hyperparameter setup of C3A for the GLUE benchmark.
+Model Hyperparameter Pets Cars DTD EuroSAT FGVC RESISC
+Both Optimizer AdamW
+LR Schedule None
+C3A Initialization Xavier Uniform
+Epochs 10
+Batch Size 64
+Base Learning Rate (C3Ab=768/12) 4E-1 4E+0 2E+0 2E+0 7E+0 2E+0
+Learning Rate (Head) 1E-2 1E-2 2E-2 8E-3 1E-2 2E-2
+Weight Decay 3E-4 5E-4 6E-5 2E-5 1E-5 2E-5
+Large Learning Rate (C3Ab=1024/16) 7E-1 4E+0 2E+0 2E+0 4E+0 3E+0
+Learning Rate (Head) 3E-3 8E-3 7E-3 2E-2 1E-1 4E-3
+Weight Decay 4E-3 1E-5 2E-4 5E-4 2E-5 9E-5
+Table A5: Hyperparameter setup of C3A for image classification tasks.
+Hyperparameter Common Math Code
+Optimizer AdamW
+LR Scheduler Cosine
+Batch Size 16 128 128
+Warmup Ratio 0.05 0.03 0.03
+Epoch 3 2 2
+Dropout 0.05
+Learning Rate (LLaMA2-7B) 3E-1 4E-1 6E-1
+Learning Rate (LLaMA3-8B) 3E-1 5E-1 6E-1
+Table A6: Hyperparameter setup of C3A for instruction tuning.